DFA

The Address utility for the AMIGA and OS 2.0 Version 2.2

by Dirk Federlein

DFA \begin{matrix}c \\ ◯ \end{matrix} 1992–1994, Dirk Federlein DFA is an SASG product gui created using Gadtoolsbox (37.300) \begin{matrix}c \\ ◯ \end{matrix} 1991–93 Jan van den Baard. arexx interface created using Arexxbox (V1.12) \begin{matrix}c \\ ◯ \end{matrix} 1992–93 Michael Balzer. The MagicWB icons are \begin{matrix}c \\ ◯ \end{matrix} Copyright 1993–94 Martin Huttenloher Installer and Installer project icon \begin{matrix}c \\ ◯ \end{matrix} Copyright 1991-93 Commodore-Amiga, Inc. All Rights Reserved. Reproduced and distributed under license from Commodore. INSTALLER SOFTWARE IS PROVIDED "AS-IS" AND SUBJECT TO CHANGE; NO WAR-RANTIES ARE MADE. ALL USE IS AT YOUR OWN RISK. NO LIABILITY OR RESPON-SIBILITY IS ASSUMED. Registered trademarks are not marked separately. Therefore absence of a trademark does not imply they are free.

Chapter 1: Welcome! Page 1

1 Welcome!

Welcome to the documentation of DFA! You did really manage it and forced yourself to read this manual! Please, do not stop here, but continue reading – it will be to your advantage!

Only by reading this manual carefully, can learn everything about the many different features of DFA! Some of your questions – if not all of them – which you might have at this time, will be answered here.

Please remember that I put a lot of time and effort (‘blood, sweat and tears’) into this manual — please show your acknowledgement by reading this documentation (of course, you may feel free to register)!

Dirk Federlein

Page 2 DFA V2.1

2 Features

DFA supplies many outstanding features, some of them I’d like to mention here:

• OS2.0 "Look and Feel"

• Locale support (Workbench 2.1 and better needed)

• OS3.0 is supported, if available

• Commodity support, i.e. DFA can be reached by hotkey or the Exchange utility of the Workbench

• All operations in DFA are accessable via keyboard

• The Arexx interface contains more than 30 commands

• All windows of DFA are font sensitive

• The DFA-Editor main window is font sensitive and resizable

• The DFA-Editor main window is an application window

• The program has been split into the editor, server and preferences parts; by this the available memory is optimally used and DFA often needs even less memory than in version 1.2x

• Program parts used by more than one part of DFA have been exported into a shared library,

  the ‘dfa.library’

• Button row, which contains the most important functions

• Support of external files

• Up to three EMail addresses per entry

• Dial function (supports OwnDevUnit) recognizes BUSY, NO DIALTONE, etc.

• The addresses can be separated into groups

• The groups can be named (incl. shortcut)

• New fields state and c/o

• Network support

• Application icon (can be toggled)

• It is possible to use DFA as default tool

• New "magic" icons (8 colors) by Martin Huttenloher

• Icon generation for the address files can be switched off

• The auto save feature can be configured

• The format of the address listview can be changed

• The display part in the main window of the DFA-Editor can be configured

• Addresses can be sorted by any field of the address

• The fonts for the DFA-Editor and the DFA-Preferences can be changed

• The utility NewEdit is supported

• The internal memory management was rewritten from scratch with the consequence that the addresses now use much less Ram than in version 1.2x

DFA — the address utility for the AMIGA Chapter 2: Features Page 3

• The usage of memory pools decreases memory fragmentation

• The load and save routines have been rewritten from scratch and are much faster now

3 Installation

I strongly recommend use of the supplied install program to install DFA. It extracts the needed file from their archives and copies them to the correct places on your harddisk. It sets all needed tooltypes as well.

After the installation procedure has been completed, there exist the following files, or according to your installation at least some of them, on your harddisk:

• ‘DFA’, ‘DFA.info’

The DFA-Server program; it supplies the Arexx–Port and manages the Application icon. Putting it into the ‘WBStartup’ drawer is a good idea.

• ‘DFAEditor’, ‘DFAEditor.info’

The graphical user interface of DFA. If you want to start this program from the Shell, please copy it into a drawer in your search path, otherwise in any directory you like, e.g. ‘DFA:c’, which is the default.

• ‘DFAPrefs’, ‘DFAPrefs.info’

The preferences program for DFA. All settings can be done from within this program. A good place for it is the ‘SYS:Prefs’ or the ‘DFA:Prefs’ directory.

• ‘dfa.library’

Must be in ‘libs:’ or in ‘DFA:libs’. This library is needed by all other parts of DFA, i.e. the DFAServer, the DFAEditor and the DFAPrefs program.

• ‘Registration’, ‘Registration.info’

This is the new registration program of the SASG. This program makes it possible to register or update in a very comfortable and fast way. Furthermore it contains detailed information about the SASG and the advantages that you may take out of it.

• ‘DFA.guide’

The amigaguide documentation of DFA. To read the file you need e.g. the AmigaGuide or the Multiview program.

• ‘DFA.dvi’

The manual in the DVI format. To look at it or print it you need TEX (e.g. PasTEX \begin{matrix}c \\ ◯ \end{matrix} Georg Heßmann).

• ‘DFAEditor.guide’

The online help for the DFA-Editor; it can be found in ‘Help:<language>’ or in the drawer you selected during the installing procedure.

• ‘DFAPrefs.guide’

The online help for the DFA-Preferences program; it can be found in ‘Help:<language>’ or in the drawer you selected during the installing procedure.

DFA — the address utility for the AMIGA Chapter 3: Installation Page 5

• ‘dfa.key’

If you’ve got a registered version of DFA, you can find the so-called keyfile ‘dfa.key’ in the ‘s’ drawer of the distribution. To give DFA the chance to find the keyfile you either have to copy it to the ‘s:’ drawer of your boot partition (default) or, if you prefer another directory for your keyfile, you will have to set the environment variable KEYPATH to this drawer.

Example:

     setenv KEYPATH ENV:DFA

Usually the install program has already done this for you.

• ‘s/adr.file’

The default address file. It contains the address of the program author. This can be useful for bug reports and — most of all — when you want to register.

• ‘rexx/#?.ced’

Arexx scripts for the CED (\begin{matrix}c \\ ◯ \end{matrix} 1987–1993 CygnusSoft Software) working together with DFA. The explanation of each Arexx script can be found in the scripts.

• ‘rexx/#?.dfa’

Arexx scripts that can be started directly out of the DFA-Editor (via the function keys).

• ‘rexx/#?.rexx’

Arexx scripts that can be be started out of the shell (using the rx command.

• ‘locale/Catalogs/<language>/#?.catalog’

In each ‘language’ directory (e.g. ‘deutsch’ you can find the catalog files that are needed for the locale support (Workbench 2.1 and better needed). Usually the installer script has already copied the needed files to the correct drawers.

4 Quick Start

After you have installed DFA (see Chapter 3 [Installation], page 4, for more information), the builtin default settings are used. These settings already allow you to use DFA and take a lot of advantage out of it. Please notice that the huge amount of settings, which can be tuned in the DFA-Preferences program, can only be saved in the registered version (see Chapter 9 [The DFA-Preferences program], page 58 and Chapter 11 [Registration], page 73 for details)!

If you copied the DFA-Server (filename ‘DFA’) into the ‘WBStartup’ drawer, as it is recommended by the install program, DFA will be started, each time the computer starts up; futhermore, an application icon for DFA is supplied on the Workbench screen.

This behaviour of DFA can be changed as described in Section 7.1 [Tooltypes], page 10 and Chapter 9 [The DFA-Preferences program], page 58.

As DFA tries to follow the "User Interface Style Guide" as closely as possible, it should be no problem (at least for advanced users) to figure out all the functions of DFA according to "trial and error". If you don’t want any bad (?) surprises, just continue reading… Chapter 5: The concept of DFA Page 7

5 The concept of DFA

Since version 2.0 of DFA has been finished, it is no longer a single program, but consists of several parts: The DFA-Server (see Chapter 7 [DFAServer], page 10), the DFA-Editor (see Chapter 8 [The DFA-Editor program], page 41) and the DFA-Preferences program (see Chapter 9 [The DFA-Preferences program], page 58).

Certain functions that are used by more than one of these programs have been exported into a socalled "shared library", the dfa.library.

By doing so it is possible now that DFA often uses less Ram than in former versions of DFA, even though it contains many enhancements! If DFA would still have been one single program it would have been a size of about 250 KB and nearly noone would have copied it into the

‘WBStartup’ drawer.

In spite of splitting the program into several parts, I tried to keep the known handling of DFA: One should think he works with one single program. To do so, the DFA-Server calls the DFA-Editor after the user hits the hotkey; the DFA-Preferences program on the other side can be reached directly out of the DFA-Editor by selecting the corresponding menu item.

This new concept will although be useful for people that start DFA on demand only and therefore don’t need the Arexx and the Commodity part of DFA; these people may start the DFA-Editor directly (from the Workbench or Shell) and save the memory the DFA-Server would have taken (about 35 KB).

6 Usage of DFA

You can control DFA by mouse as usual as well as completely by keyboard.

Some notes to the usage of the keyboard:

• Button Gadgets All Gadgets have a corresponding text. In this text you’ll find one character underlined. This character is the "shortcut" to the gadget. To use this shortcut, just push the corresponding key. You don’t have to push any qualifier like CONTROL or ALT. To simple buttons it doesn’t matter if you push the "normal" or the capital letter. To other gadgets it can make a difference (see below for details).

• Listviews Listviews are used in the DFA-Editor as well as in the DFA-Preferences program. There are two different types of listviews: Listviews in which you can select entries (raised border), e.g. the address listview in the DFA-Editor main window, and listviews that only display information (recessed border), e.g. the listview in the DFA-Editor full window. If not mentioned otherwise, all listviews can be used as follows: Input listview:

Shortcut The next entry becomes the active one, the list scrolls accordingly.

Shift Shortcut The previous entry becomes the active one, the list scrolls accordingly.

Read only listview:

Shortcut The list scrolls up.

Shift Shortcut The list scrolls down.

Sometimes the listviews can be controlled by the cursor keys or it is possible to scroll page by page. If this is the case, the enhanced usage of that particular listview is described seperately on that place.

• Cycle Gadgets In front of a cycle gadget there is short text, which describes the function of the gadget. Again there is one character underlined. However, here it does matter if you use the capital letter or not. The "normal" letter cycles foreward, the capital letter backward.

• Radio Buttons Only one of the buttons that belong together can be activated at once. You can select an entry by pushing the proper shortcut. The case of the shortcut is not important.

Chapter 6: Usage of DFA Page 9

• Check Boxes A check box can be selected (shown by a tick) or deselected ("blank"). You can switch the current state by pushing the corresponding hotkey.

• String Gadgets The string gadgets have the same functions as the "original" string gadgets. However, I added some features, which should make them easier to use: − Hit RETURN to end the input to the current gadget and go to the next string or integer

gadget. You may use the ENTER key alternatively. − Hit SHIFTRETURN to end the input to the current gadget and go to the previous gadget. − CURSOR UP or CURSOR DOWN brings the cursor to the previous or next string gadget

respectively. − Hit ESCAPE to cancel the input completely. The window which contains the gadget is

closed and all previous input is discarded. − If the cursor stands within a string gadget, the shortcuts can be reached by hitting

RIGHT AMIGA–<letter> and SHIFT. If you don’t hit SHIFT additionally, the keyboard shortcut is used either by the internal string gadget edit features (like RIGHT AMIGA X or RIGHT AMIGA Q) or by additional tools like NewList.

You activate an string gadget by pushing the according shortcut. Usually there is no difference if you hit the upper case or the lower case character, as long as there is no button gadget behind the string gadget. This button gadget that contains a small symbol instead of text, activates a requester, which allows a more comfortable input in the string gadget infront of it. If there is such a button, you activate the string gadget with the lower case key and the button gadget behind the string gadget with the upper case key.

• Text Gadgets Usually text gadgets don’t have a shortcut, as they only display text, but don’t allow any input. Some text gadgets however have a small button gadget behind them (cf. DFA-Preferences: Font–Preferences) After you have hit this gadget a requester pops up and you may select an entry which is displayed in the text gadget as soon as you’ve completed the requester. Therefore the gadget shortcut doesn’t refer to the text gadget itself, but to the button gadget behind it.

• Online Help Since version 1.2 has been finished, DFA supports an Online Help feature. To invoke it, just hit the Help key and the help window pops up. Please remember that the ‘GUIDEFILE’ tooltype of the DFA-Editor and the DFA-Preferences program have to be set correctly. The installer program usually has already done this for you.

7 The DFA-Server program

The DFA-Server (filename: ‘DFA’) supplies the commodity features (hotkey, access by the Exchange program) and contains the complete Arexx interface; furthermore it manages the application icon.

7.1 Tooltypes

If you click on the icon of the DFA-Server program once and activate the information program from the Workbench menu you may change the following tooltypes:

• ADDRESSFILE

If you use this tooltype, you can change the name of the address file that will be loaded by DFA during startup (e.g. ‘ADDRESSFILE=MyAddresses.dfa’). Please notice that in this case the file you set in the DFA-Preferences program is not used.

• CX_POPKEY

The preset hotkey is LCOMMAND F5, i.e. you can pop up the DFA-Editor by pushing Left

Amiga F5 (cf. ‘CX_POPUP’).

You can change this hotkey using the following keywords:

LSHIFT left shift key RSHIFT right shift key CAPSLOCK caps lock key CONTROL control key LALT left alt key RALT right alt key LCOMMAND left AMIGA key RCOMMAND right AMIGA key

NUMERICPAD numeric pad

MIDBUTTON middle mouse button RBUTTON right mouse button LEFTBUTTON left mouse button F1-F10 function keys 1 to 10 UP cursor up DOWN cursor down LEFT cursor left RIGHT cursor right HELP help key DEL del key RETURN return key ENTER enter key on the numeric pad; notice that you have

to set NUMERICPAD as well!

Chapter 7: The DFA-Server program Page 11

BACKSPACE backspace key ESC escape key SPACE space key COMMA comma key UPSTROKE upstroke key

     A..Z, a..z  "usual" keys
     0..9, etc.

You can "mix" these keys, for example:

 − LSHIFT CONTROL F1
 − LALT CONTROL D
 − CONTROL NUMERICPAD ENTER
 − RCOMMAND RSHIFT A
• CX_POPUP

If you set CX_POPUP=YES, you make the DFA-Server calling the DFA-Editor immediately after startup. Set CX_POPUP=NO to keep DFA "quiet" until you pop up the DFA-Editor by hitting the hotkey or clicking on the application icon or by starting the DFA-Editor directly from the Workbench or the Shell.

• CX_PRIORITY

With this tool type you can change the Priority of the DFA commodity part. The default setting of CX_Priority=0 usually doesn’t need to be changed. For more information look up your manual to the AMIGA–OS.

• DFAEDITOR

This tooltype contains the filename (incl. the full path) that is used by the DFA-Server to call the DFA-Editor. The install program has usually set this tooltype according to your wishes, you only have to change anything if you copied the DFA-Editor manually to another place on your harddisk.

• DONOTWAIT

This is a system tooltype, which forces the Workbench not to wait for DFA ending. You should not remove this tool type.

• PORTNAME

DFA uses ‘DFA’ for the portname. If there is another program with the same portname, DFA attaches numbers (‘DFA.1’, ‘DFA.2’, etc.) until a definite name is found. However, if you want your "own" portname, you can set it using the tool type PORTNAME (e.g. PORTNAME=OtherDFAPortname). Please remember that you have to adapt all example arexx scripts, if you change the portname!

7.2 CLI–Parameter

If you start DFA from Workbench the tooltypes described above will be used. When you start DFA from Shell, these tooltypes are used as well, as long as the corresponding .info-File is present. You can give the following parameters to DFA that overwrite the previous settings (within the .info file). The possible arguments are described below:

ADDRESSFILE/K, CX_POPKEY/K, CX_PRIORITY/N, CX_POPUP/K, PORTNAME/K, DFAEDITOR/K

This means, you may…

. . . set the hotkey to Left Amiga F1 by ‘DFA CX_POPUP="LCOMMAND F1"’
. . . set the Arexx port to ‘MyArexxPort’ by ‘DFA PORTNAME=MyArexxPort’
. . . load the alternate address file ‘MyAddresses.dfa’ by ‘DFA ADDRESSFILE=MyAddresses.dfa’

7.3 Commodity

By pushing the hotkey (see Section 7.1 [Tooltypes], page 10) you may envoke the DFA-Editor. You’ll get the same effect, if you start the ‘Exchange’ program of the Workbench and select

‘Show Interface’.

Using the ‘Exchange’ program you can although stop DFA (hit ‘Remove’) or to make it inactive (hit ‘Inactive’) or active respectively.

7.4 Application–Icon

The application icon supplies two functions:

• A double click on it activates the DFA-Editor.

• If you let the icon of an address file "fall" on the application icon, the DFA-Editor is started and the according address file is loaded.

7.5 Arexx

7.5.1 Important — new features since version 1.2x

The Arexx interface has been rewritten from scratch for DFA 2.0. Some commands have been added…

• ATTEMPTLOCK

• CHANGEGROUPS

• FREELOCK

• NEWFILE … some old commands have been changed or improved regarding their syntax:

DFA — the address utility for the AMIGA • APPEND

• EDIT

• KILL

• LOAD

• NEW

• SAVEAS

• SEARCH

Most of all, however, I’ve changed the order of the fields within an address item, which is returned by several commands like ‘FIRST’, ‘NEXT’, ‘GETCURRENT’, etc.

The reason for doing so is that I thought it would be a good idea to have the same order of the fields in the Arexx interface as in the DFA-Editor edit window, as this makes it a lot easier to remember which meaning the different numbers of an address field have (cf. Section 7.5.3 [Format of the address field], page 15).

7.5.2 Basics

DFA (i.e. the DFA-Server) usually accepts Arexx commands at any time, even if the DFA-Editor and/or the DFA-Preferences program is running at the same time. It is possible however that the execution of an Arexx command will be delayed, if someone for example is just editing an address within the Edit window of the DFA-Editor. The reason for this savety protocol can be explained easily: As the addresses are kept in memory only once, all DFA related programs have access to the same address list. By the protocol described above I want to make sure that no part of the DFA package tries to use addresses that another part of DFA has already deleted (this would sooner or later cause the machine to crash).

When I talk about executing Arexx commands, the execution of an Arexx script that contains DFA specific Arexx commands, is meant.

The basics of the Arexx programming language can be found in the Arexx manual that comes along with AMIGA–OS 2.0 or better or in the "Arexx User’s Reference Manual"^{1}.

The Arexx port, that has to be used to address DFA is ‘DFA’, as long as you did not change this name by the tooltype PORTNAME, as described in Section 7.1 [Tooltypes], page 10. This should only be done, if you have another program running that uses the portname ‘DFA’ as well and you don’t want to use DFA’s feature to choose a clear portname.

A further remark on the Load, Append, and Save as commands:

1 Arexx User’s Reference Manual, William S. Hawes, P.O. Box 308, Maynard, MA 01754 If you give a filename, you should always use the complete pathname, as otherwise it is very likely that the file can’t be found or is written to a place you don’t expect. Refering to the home directory of DFA doesn’t make much sense, as DFA is often placed in another drawer than the address files and furthermore as DFA remembers the recently loaded filename and uses it for loading and saving.

Beyond the item Syntax the possible parameters are listed. The description of the parameters follows the known style given by Commodore, here is its meaning:

/S – Switch This is considered a boolean variable. If this option is present, the corresponding option is enabled, otherwise it keeps disabled.

/K – Keyword.

This parameter must be given in the format ‘keyword=<setting>’; e.g. the ‘PRINT’

command can get FORM/K either as FORM=ALL or FORM=SELECTED or FORM=ACTIVE.

The equal sign = may be left out.

/N – Number.

This means the parameter is considered a decimal integer,

/T – Toggle.

This is similar to the switch (/S) modifier, but one of the keywords yes, on, no or off is expected (not case sensitive). According to the keyword you provide, the option is switched on or off respectively. Please notice that you may not use the equal sign (=) between the parameter and the modifier. The following examples may give you an idea, how this switch can be used:

Examples:

         gui input off output on
         gui input yes output ON
         gui input NO output off

Important: Starting with version 2.2 of DFA the behaviour of this /T switch changed a bit, so please adopt your Arexx scripts accordingly.

/A – Always.

This modifier means that this option is required. It must appear in the command line.

The exact syntax for every Arexx command can be found below, see Section 7.5.4 [Arexx commands], page 16, for details.

Important:

The Arexx interface of DFA gives you a powerful instrument to manage the stored addresses. A wider range of functionality however implements more ways of loosing your data! This could happen not only by a malfunction of DFA but also by an inproper used function or parameter. Therefore I encourage you to make regular copies of your databases to keep the potential damage as small as possible.

7.5.3 Format of the Address Field

If there is listed ADDRESS/M in the return code section, this means that the corresponding command returns the wanted address if it has been executed successfully. The address can be found either in the variable given together with the Arexx keyword STEM or VAR or in the default variable RESULTS.

Example:

     /* Outputs the first address */
     ADDRESS "DFA"
     OPTIONS RESULTS
     /* Please notice the full stop behind 'TEST' ! */
     FIRST STEM TEST.
     SAY TEST.ADDRESS.2
     EXIT

This short example outputs the name of the first entry of the address list.

The STEM keyword returns the address in the form of a list, i.e. the address items appear

in ‘TEST.ADDRESS.0’ through ‘TEST.ADDRESS.25’. In ‘TEST.ADDRESS.COUNT’ you can find the

number of entries of the address array. The VAR keyword lets DFA return the desired address as well, but the whole address is stored in one single variable, separated by spaces. Please consult your Arexx manual for further details.

The address fields correspond to the following "numbers":

0 Salutation

1 First 2 Name 3 Co 4 Street 5 Zip 6 City 7 State 8 Country 9 Birthday 10 Phone 11 Fax 12 Email1 13 Email2 14 Email3 15 Comment 16 Group1 17 Group2 18 Group3 19 Group4 20 Group5 21 Group6 22 Group7 23 Group8 24 Selected 25 External

7.5.4 Arexx commands

7.5.4.1 The ABOUT Command

‘Syntax:’ ABOUT
‘Function:’

The About window of DFA is opened DFA — the address utility for the AMIGA

7.5.4.2 The APPEND Command

‘Syntax:’ APPEND [FILENAME] <string> [FORCE] [PROMPT]

‘Argument template:’

FILENAME, FORCE/S, PROMPT/S

‘Function:’

New addresses are appended to the already existing address database. If the already existing addresses have been changed, the new addresses are not appended, as long as you don’t supply the keyword FORCE additionally. It is possible to use the PROMPT parameter, which will supply a file requester to input the desired filename of the address database to append.

‘Argument description:’

• FILENAME <string> The filename of the address file to be appended.

• FORCE Forces appending the address file.

• PROMPT Opens a file requester that allows the selection of the address file to be appended.

‘Results:’

The following error codes can be returned in RC2:

• RXERR NOFILENAME

• RXERR APPENDFAILED

• RXERR ENVCHANGED

• RXERR NONETWORKAPPEND

• RXERR MODIFIED

7.5.4.3 The ATTEMPTLOCK Command

‘Syntax:’ ATTEMPTLOCK [<retries>] [<delay>]

‘Argument template:’

RETRIES/N, DELAY/N

‘Function:’

If you are in network mode (see Section 9.4.5 [Networking], page 61), you have to get write access to the address file, before you may modify anything. To do so you have to use the ‘ATTEMPTLOCK’ command. If someone else in the Lan keeps the write access to the desired address file, this commands returns an error code in RC and RC2 (see Section 7.5.4.10 [FREELOCK], page 23).

‘Argument description:’

DFA — the address utility for the AMIGA • RETRIES

Maximal number of retries, if the attempt to get write access did not succeed at once.

• DELAY

Number of seconds that shall pass between two retries.

‘Results:’

The following error codes can be returned in RC2:

• RXERR ENVCHANGED

7.5.4.4 The CHANGEGROUPS Command

‘Syntax:’ CHANGEGROUPS [GROUP1] [GROUP2] [GROUP3] [GROUP4] [GROUP5]

[GROUP6] [GROUP7] [GROUP8] [PROMPT]

‘Argument template:’

GROUP1/S, GROUP2/S, GROUP3/S, GROUP4/S, GROUP5/S, GROUP6/S, GROUP7/S, GROUP8/S, PROMPT/S

‘Function:’

Changes the active group selection. Commands like ‘PRINT’ or ‘SEARCH’ apply only to the currently active groups. To reach all addresses in all groups you have to activate them using this command before.

‘Argument description:’

• GROUP1

• GROUP2

• GROUP3

• GROUP4

• GROUP5

• GROUP6

• GROUP7

• GROUP8

• PROMPT

Opens a window, in which the user may select the desired groups.

‘Results:’

The following error codes can be returned in RC2:

• RXERR NOFILENAME

• RXERR NOMEM

• RXERR ENVCHANGED

• RXERR MODIFIED

‘Example:’

DFA — the address utility for the AMIGA

               /* Activates groups 1 through 6 and */
               /* deactivates group 7 and 8        */
               ADDRESS 'DFA'
               CHANGEGROUPS GROUP1 GROUP2 GROUP3 GROUP4 GROUP5 GROUP6

7.5.4.5 The CLEARALL Command

‘Syntax:’ CLEARALL

‘Function:’

All items of the address list are deselected.

‘Results:’

The following error codes can be returned in RC2:

• RXERR ENVCHANGED

• RXERR NONETWORKCLEAR

‘Note:’

Only addresses that correspond to the current group selection are deselected!

7.5.4.6 The DESELECT Command

‘Syntax:’ DESELECT

‘Function:’

The current entry (it one exists) is deselected.

‘Results:’

The following error codes can be returned in RC2:

• RXERR NOCURRENT

• RXERR ENVCHANGED

7.5.4.7 The DIAL Command

‘Syntax:’ DIAL
‘Function:’

The current address (if one exists) is dialed up.

‘Results:’

The following error codes can be returned in RC2:

DFA — the address utility for the AMIGA • RXERR SERNOCARRIER

• RXERR SERNODIALTONE

• RXERR SERRING

• RXERR SERBUSY

• RXERR SEROK

• RXERR SERERROR

• RXERR SERCONNECT

• RXERR SERVOICE

• RXERR SERUNKNOWN

• RXERR NOSERIALDEV

• RXERR NOPHONENUM

• RXERR ENVCHANGED

7.5.4.8 The EDIT Command

‘Syntax:’ EDIT [SALUTATION <string>] [FIRST <string>] [NAME <string>] [STREET <string>] [CO <string>] [ZIP <string>] [CITY <string>] [STATE <string>] [COUN-TRY <string>] [BIRTHDAY <string>] [PHONE <string>] [FAX <string>] [EMAIL1 <string>] [EMAIL2 <string>] [EMAIL3 <string>] [COMMENT <string>] [GROUP1]

[GROUP2] [GROUP3] [GROUP4] [GROUP5] [GROUP6] [GROUP7] [GROUP8]

[SELECT] [EXTERNAL <string>]

‘Argument template:’

SALUTATION/K, FIRST/K, NAME/K, STREET/K, CO/K, ZIP/K, CITY/K, STATE/K, COUNTRY/K, BIRTHDAY/K, PHONE/K, FAX/K, EMAIL1/K, EMAIL2/K, EMAIL3/K, COMMENT/K, GROUP1/T, GROUP2/T, GROUP3/T, GROUP4/T, GROUP5/T, GROUP6/T, GROUP7/T, GROUP8/T, SELECT/T, EXTERNAL/K

‘Function:’

The current address (if one exists) is modified in the given fields. The former contents of these fields is deleted. Fields that are not given as parameter are not changed and keep their former contents. Please notice that parameters that contain spaces have to be set into quotes, for example:

     EDIT 'COMMENT="A test that contains spaces"'

Please remember to use ' for the outer and " for the inner quotes! Since version 2.2 of DFA has been completed, the behavior of the /T modifiers has slightly changed. If you want to switch a /T parameter on, you have to supply a ‘ON’ or ‘YES’ just behind the according parameter. To switch a parameter off, use the keywords ‘OFF’ or ‘NO’, e.g.

       EDIT GROUP1 ON GROUP2 OFF GROUP8 ON SELECT ON
‘Argument description:’

• SALUTATION <string> Salutation

• FIRST <string> First name

• NAME <string> Name

• STREET <string> Street

• CO <string> c/o

• ZIP <string> Postal code

• CITY <string> City

• STATE <string> State

• COUNTRY <string> Country

• BIRTHDAY <string> Birthday

• PHONE <string> Phone number

• FAX <string> Telefax number

• EMAIL1 <string> Email address No. 1

• EMAIL2 <string> Email address No. 2

• EMAIL3 <string> Email address No. 3

• COMMENT <string> Comment

• GROUP1 Group No. 1

• GROUP2 Group No. 2

DFA — the address utility for the AMIGA • GROUP3

Group No. 3

• GROUP4 Group No. 4

• GROUP5 Group No. 5

• GROUP6 Group No. 6

• GROUP7 Group No. 7

• GROUP8 Group No. 8

• SELECT Selection state, i.e. ‘ON’ or ‘OFF’.

• EXTERNAL <string>

The filename of the external file.

‘Results:’

The following error codes can be returned in RC2:

• RXERR NODIR

• RXERR NOFILE

• RXERR NOCURRENT

• RXERR ENVCHANGED

• RXERR NONETWORKEDIT

7.5.4.9 The FIRST Command

‘Syntax:’ FIRST [VAR <name>] [STEM <name>]

‘Argument template:’

VAR/K, STEM/K

‘Function:’

The first address of the address list becomes the current one and is returned in the ADDRESS variable.

‘Argument description:’

• VAR <name>

• STEM <name>

‘Results:’

• .ADDRESS.COUNT

• .ADDRESS.x

DFA — the address utility for the AMIGA The following error codes can be returned in RC2:

• RXERR NOFIRST

• RXERR ENVCHANGED

7.5.4.10 The FREELOCK Command

‘Syntax:’ FREELOCK

‘Function:’

If you’ve got write access using the ATTEMPTLOCK command (cf. Section 7.5.4.3 [ATTEMPTLOCK], page 17), you should free the locked file as soon as possible using the FREELOCK command, to give other members of the Lan the chance to access this address file.

‘Results:’

The following error codes can be returned in RC2:

• RXERR NOSAVE

• RXERR ENVCHANGED

‘Note:’

Makes sense in network mode only!

7.5.4.11 The GETCURRENT Command

‘Syntax:’ GETCURRENT [VAR <name>] [STEM <name>]

‘Argument template:’

VAR/K, STEM/K

‘Function:’

The current entry (if there exists one) is returned.

‘Argument description:’

• VAR <name>

• STEM <name>

‘Results:’

• .ADDRESS.COUNT

• .ADDRESS.x The following error codes can be returned in RC2:

• RXERR NOCURRENT

• RXERR ENVCHANGED DFA — the address utility for the AMIGA

7.5.4.12 The GUI Command

‘Syntax:’ GUI [<input>] [<output>]

‘Argument template:’

INPUT/T, OUTPUT/T

‘Function:’

Using this command you may lock or free the DFAEditor’s output and input respectively. After you’ve locked the input of the DFAEditor using ‘GUI INPUT OFF’, it is no longer possible to do any input to the DFAEditor, i.e. you cannot select any gadgets or menus. If you’ve used ‘GUI OUTPUT OFF’ to lock the output of the DFAEditor, the address and teh panel listview of the DFAEditor main window are no longer updated, when you more through the address list using ‘NEXT’ or ‘PREV’ for example. Not before you’ve used the command ‘GUI OUTPUT ON’, the displayed is "activated" again.

‘Argument description:’

• INPUT Disables or enables user input into the DFAEditor.

• OUTPUT

Disables the refreshing of the DFAEditor’s address and panel listview.

‘Results:’

The following error codes can be returned in RC2:

• RXERR SYNTAX

‘Note:’

Please make sure that you enable input and/or output, before your Arexx script ends, if you’ve formerly disabled them.

7.5.4.13 The ICONIFY Command

‘Syntax:’ ICONIFY

‘Argument template:’
‘Function:’

The DFA-Editor is stopped, if it is currently running. The DFA-Server, however, keeps running, so you still can reach the DFA-Editor by pushing the hotkey.

‘Results:’

The following error codes can be returned in RC2:

• RXERR NOICONIFY

7.5.4.14 The KILL Command

‘Syntax:’ KILL [CURRENT] [SELECTED]

‘Argument template:’

CURRENT/S, SELECTED/S

‘Function:’

The current entry (CURRENT) or all selected (SELECTED) addresses are deleted! If you use the KILL command without a parameter, it has the same meaning as if you had given the CURRENT parameter, i.e. the current address (as long as there is one) is deleted.

‘Argument description:’

• CURRENT The current address will be deleted. Can be left out.

• SELECTED

All selected addresses will be deleted.

‘Results:’

The following error codes can be returned in RC2:

• RXERR ENVCHANGED

• RXERR NONETWORKKILL

‘Note:’

No safety requester will pop up! The addresses are deleted even if they have been modified before!

7.5.4.15 The LAST Command

‘Syntax:’ LAST [VAR <name>] [STEM <name>]

‘Argument template:’

VAR/K, STEM/K

‘Function:’

The last address becomes the current one and is returned in ADDRESS.

‘Argument description:’

• VAR <name>

• STEM <name>

‘Results:’

• .ADDRESS.COUNT

• .ADDRESS.x

‘Results:’

The following error codes can be returned in RC2:

• RXERR ENVCHANGED

DFA — the address utility for the AMIGA

7.5.4.16 The LOAD Command

‘Syntax:’ LOAD [FILENAME] <string> [FORCE] [PROMPT]

‘Argument template:’

FILENAME, FORCE/S, PROMPT/S

‘Function:’

The given address file is loaded, if the previously loaded one had not been changed. You may force the loading of the new addresses by giving the FORCE parameter. Please notice that using the FORCE option throws away any changes you applied to the former loaded addresses. Instead of giving a filename it is possible to use the PROMPT parameter, which will supply a file requester to input the desired filename.

‘Argument description:’

• FILENAME <string> Address file to be loaded.

• FORCE Forces loading a new address file.

• PROMPT

Supplies a file requester to select the desired filename.

‘Results:’

The following error codes can be returned in RC2:

• RXERR NOLOAD

• RXERR ENVCHANGED

• RXERR MODIFIED

• RXERR NOFILENAME

7.5.4.17 The NEW Command

‘Syntax:’ NEW [SALUTATION <string>] [FIRST <string>] [NAME <string>] [STREET <string>] [CO <string>] [ZIP <string>] [CITY <string>] [STATE <string>] [COUN-TRY <string>] [BIRTHDAY <string>] [PHONE <string>] [FAX <string>] [EMAIL1 <string>] [EMAIL2 <string>] [EMAIL3 <string>] [COMMENT <string>] [GROUP1]

[GROUP2] [GROUP3] [GROUP4] [GROUP5] [GROUP6] [GROUP7] [GROUP8]

[SELECT] [EXTERNAL <string>]

‘Argument template:’

SALUTATION/K, FIRST/K, NAME/K, STREET/K, CO/K, ZIP/K, CITY/K, STATE/K, COUNTRY/K, BIRTHDAY/K, PHONE/K, FAX/K, EMAIL1/K, EMAIL2/K, EMAIL3/K, COMMENT/K, GROUP1/S, GROUP2/S, GROUP3/S, GROUP4/S, GROUP5/S, GROUP6/S, GROUP7/S, GROUP8/S, SELECT/S, EXTERNAL/K

‘Function:’

A new address is created using the provided data.

‘Argument description:’

• SALUTATION <string> Salutation

• FIRST <string> First name

• NAME <string> Name

• STREET <string> Street

• CO <string> c/o

• ZIP <string> Postal code

• CITY <string> City

• STATE <string> State

• COUNTRY <string> Country

• BIRTHDAY <string> Birthday

• PHONE <string> Phone number

• FAX <string> Telefax number

• EMAIL1 <string> EMail address No. 1

• EMAIL2 <string> EMail address No. 2

• EMAIL3 <string> EMail address No. 3

• COMMENT <string> Comment

• GROUP1 Group No. 1

DFA — the address utility for the AMIGA • GROUP2

Group No. 2

• GROUP3 Group No. 3

• GROUP4 Group No. 4

• GROUP5 Group No. 5

• GROUP6 Group No. 6

• GROUP7 Group No. 7

• GROUP8 Group No. 8

• SELECT Selection state, i.e. ‘0’ or ‘1’.

• EXTERNAL <string> The filename of the external file.

7.5.4.18 The NEWFILE Command

‘Syntax:’ NEWFILE [<filename>] [FORCE] [PROMPT]

‘Argument template:’

FILENAME, FORCE/S, PROMPT/S

‘Function:’

This command makes it possible to create a new address file. The command fails, if the previously loaded address file has been modifed. This behaviour can be changed by giving the FORCE option additionally. Using the parameter PROMPT enables a file requester to input the new filename.

‘Argument description:’

• FILENAME New filename.

• FORCE Forces the creation of the new file, even if the old address file has been modified.

• PROMPT

Supplies a file requester to input the new filename.

‘Results:’

The following error codes can be returned in RC2:

DFA — the address utility for the AMIGA • RXERR NODIR

• RXERR NOEXTERNAL

• RXERR ENVCHANGED

• RXERR NONETWORKNEW

7.5.4.19 The NEXT Command

‘Syntax:’ NEXT [VAR <name>] [STEM <name>]

‘Argument template:’

VAR/K, STEM/K

‘Function:’

The next address becomes the current one and is returned.

‘Argument description:’

• VAR <name>

• STEM <name>

‘Results:’

• .ADDRESS.COUNT

• .ADDRESS.x The following error codes can be returned in RC2:

• RXERR NONEXT

• RXERR ENVCHANGED

7.5.4.20 The NEXTSEL Command

‘Syntax:’ NEXTSEL [VAR <name>] [STEM <name>]

‘Argument template:’

VAR/K, STEM/K

‘Function:’

The next selected address becomes the current one and is returned.

‘Argument description:’

• VAR <name>

• STEM <name>

‘Results:’

• .ADDRESS.COUNT

• .ADDRESS.x The following error codes can be returned in RC2:

• RXERR NONEXTSEL

• RXERR ENVCHANGED

DFA — the address utility for the AMIGA

7.5.4.21 The OPENEDITWINDOW Command

‘Syntax:’ OPENEDITWINDOW

‘Argument template:’

‘Function:’

The Edit window is opened and the current address is displayed, if there is one. The execution of the Arexx script is halted until the Edit window is closed.

‘Argument description:’

-

‘Results:’

The following error codes can be returned in RC2:

• RXERR CURRENT

‘Warning:’

The Edit window is opened only if there is a current address, otherwise an error code is returned (see above).

7.5.4.22 The POPUP Command

‘Syntax:’ POPUP
‘Function:’

If the DFA-Editor doesn’t run up to now, it is started. Then it opens its main window. If the DFA-Editor is already running, it comes to front.

7.5.4.23 The PREV Command

‘Syntax:’ PREV [VAR <name>] [STEM <name>]

‘Argument template:’

VAR/K, STEM/K

‘Function:’

The previous address becomes the current one and is returned.

‘Argument description:’

• VAR <name>

• STEM <name>

‘Results:’

• .ADDRESS.COUNT

• .ADDRESS.x

DFA — the address utility for the AMIGA The following error codes can be returned in RC2:

• RXERR NOPREV

• RXERR ENVCHANGED

7.5.4.24 The PREVSEL Command

‘Syntax:’ PREVSEL [VAR <name>] [STEM <name>]

‘Argument template:’

VAR/K, STEM/K

‘Function:’

The previously selected address becomes the current one and is returned.

‘Argument description:’

• VAR <name>

• STEM <name>

‘Results:’

• .ADDRESS.COUNT

• .ADDRESS.x The following error codes can be returned in RC2:

• RXERR NOPREVSEL

• RXERR ENVCHANGED

7.5.4.25 The PRINT Command

‘Syntax:’ PRINT [WHICH <string>] [FILE <string>] [NLQ] [CONDENSED] [FORMFEED]

[<form>]

‘Argument template:’

WHICH/K, FILE/K, NLQ/S, CONDENSED/S, FORMFEED/S, FORM/N

‘Function:’

Prints the desired addresses.

‘Argument description:’

• WHICH <string> WHICH can get the following keywords: ALL All addresses are printed SELECTED Only the selected addresses are printed ACTIVE Only the active entry (if there is one) is printed. DFA — the address utility for the AMIGA • FILE <string>

the FILE parameter gets a filename. To this file the addresses are written instead of printed on the printer.

• NLQ Prints the addresses in NLQ

• CONDENSED Prints the addresses in condensed mode

• FORMFEED After the last address, which has been printed, a formfeed is performed. The settings from the DFA-Editor print window are not taken!

• FORM the FORM keyword can get the numbers 0 through 4. These numbers have the same meaning as the order of the print forms in the print window, but starting with 0 instead of 1.

0: Short address list 1: Long address list 2: Telephone list 3: mailing labels, [70mm x 46mm] 4: address cards

‘Results:’

The following error codes can be returned in RC2: • RXERR SYNTAX

• RXERR ENVCHANGED

• RXERR OPNPRT

• RXERR CRTEXTIO

• RXERR CRTPORT

• RXERR WRITEALL

• RXERR WRITESEL

• RXERR WRITECURRENT

7.5.4.26 The SAVE Command

‘Syntax:’ SAVE
‘Function:’

The addresses are saved under their current name.

‘Results:’

The following error codes can be returned in RC2: • RXERR ENVCHANGED

• RXERR NONETWORKSAVE

DFA — the address utility for the AMIGA

7.5.4.27 The SAVEAS Command

‘Syntax:’ SAVEAS [FILENAME] <string> [FORCE] [PROMPT]

‘Argument template:’

FILENAME, FORCE/S, PROMPT/S

‘Function:’

The addresses are saved under the given name.

‘Argument description:’

• FILENAME <string> Name of the address file to save.

• FORCE Saves the addresses even if they have not been changed before.

• PROMPT Supplies a file requester, which may be used to select the name of the address file to be saved.

‘Results:’

The following error codes can be returned in RC2:

• RXERR NONETWORKSAVE

• RXERR NOSAVE

• RXERR ENVCHANGED

7.5.4.28 The SAVEASDEFAULT Command

‘Syntax:’ SAVEASDEFAULT

‘Function:’

The addresses are saved under the default name that you entered in the DFA-Preferences (default name: ‘s:adr.file’).

‘Results:’

The following error codes can be returned in RC2:

• RXERR NONETWORKSAVE

• RXERR NOSAVE

• RXERR ENVCHANGED

‘Note:’

Please use this command very carefully! If you’ve just loaded another than the default address file, the default address file is overwritten by this command!

DFA — the address utility for the AMIGA

7.5.4.29 The SEARCH Command

‘Syntax:’ SEARCH [PATTERN] <string> [IGNORECASE] [ALL] [SALUTATION] [FIRST] [NAME] [STREET] [CO] [ZIP] [CITY] [STATE] [COUNTRY] [BIRTHDAY]

[PHONE] [FAX] [EMAIL1] [EMAIL2] [EMAIL3] [COMMENT] [VAR <name>] [STEM <name>]

‘Argument template:’

PATTERN/A, IGNORECASE/S, ALL/S, SALUTATION/S, FIRST/S, NAME/S, STREET/S, CO/S, ZIP/S, CITY/S, STATE/S, COUNTRY/S, BIRTHDAY/S, PHONE/S, FAX/S, EMAIL1/S, EMAIL2/S, EMAIL3/S, COMMENT/S, VAR/K, STEM/K

‘Function:’

The given pattern is searched for. If an address fits into the pattern, it is returned. If the DFA-Editor is currently open, the address list scrolls to the corresponding place.

‘Argument description:’

• PATTERN <string> A search pattern. The known patterns from the Shell, as for example #? may (and must) be used. Please notice that partial strings are only found, if you put an ‘#?’ before and/or behind the string (see example for details)! A detailed description of possible search patterns can be found in Section 8.9 [Search], page 54.

• IGNORECASE Upper and lower case is treated as the same.

• ALL All fields are searched.

• SALUTATION Only the Salutation fields are searched through.

• FIRST Only the First fields are searched through.

• NAME Only the Name fields are searched through.

• STREET Only the Street fields are searched through.

• CO Only the c/o fields are searched through.

• ZIP Only the ZIP fields are searched through.

• CITY Only the City fields are searched through.

DFA — the address utility for the AMIGA • STATE

Only the State fields are searched through.

• COUNTRY Only the Country fields are searched through.

• BIRTHDAY Only the Birthday fields are searched through.

• PHONE Only the Phone fields are searched through.

• FAX Only the Fax fields are searched through.

• EMAIL1 Only the EMail1 fields are searched through.

• EMAIL2 Only the EMail2 fields are searched through.

• EMAIL3 Only the EMail3 fields are searched through.

• COMMENT Only the Comment fields are searched through.

• VAR <name>

• STEM <name>

‘Results:’

• .ADDRESS.COUNT

• .ADDRESS.x The following error codes can be returned in RC2:

• RXERR ENVCHANGED

• RXERR NOTFOUND

‘Example:’
   /*
    * Look for all names beginning with M.
    * Make no difference between upper and lower case.
    */
   ADDRESS 'DFA'
   SEARCH 'M#?' IGNORECASE NAME

7.5.4.30 The SEARCHNEXT Command

DFA — the address utility for the AMIGA ‘Syntax:’ SEARCHNEXT [VAR <name>] [STEM <name>]

‘Argument template:’

VAR/K, STEM/K

‘Function:’

Continue searching for the pattern given in SEARCH before. If the search succeeded, the found address is returned. If the DFA-Editor is currently running, the address listview is updated.

‘Argument description:’

• VAR <name>

• STEM <name>

‘Results:’

• .ADDRESS.COUNT

• .ADDRESS.x

‘Note:’

The following error codes can be returned in RC2:

• RXERR NOTFOUND

• RXERR ENVCHANGED

7.5.4.31 The SELALL Command

‘Syntax:’ SELALL

‘Function:’

All addresses are selected.

‘Results:’

The following error codes can be returned in RC2:

• RXERR ENVCHANGED

• RXERR NONETWORKSELECT

‘Note:’ Only addresses that match the current group selection are selected.

7.5.4.32 The SELBYNAME Command

‘Syntax:’ SELBYNAME [PATTERN] <string>

‘Argument template:’

PATTERN/A

‘Function:’

All addresses that match the given pattern are selected. Upper and lower case does not make a difference.

DFA — the address utility for the AMIGA

‘Argument description:’

• PATTERN <string> A pattern string. The description of such patterns can be found in Section 8.9 [Search], page 54.

‘Results:’

The following error codes can be returned in RC2:

• RXERR ENVCHANGED

• RXERR NONETWORKSELECT

‘Note:’

All fields of all addresses (that match the current group selection) are searched through.

7.5.4.33 The SELECT Command

‘Syntax:’ SELECT [QUERY] [VAR <name>] [STEM <name>]

‘Argument template:’

QUERY/S, VAR/K, STEM/K

‘Function:’

The current entry (if there is one) is selected. If you supply the QUERY parameter, the current address is not selected, but its selection state (0 or 1) is returned.

‘Argument description:’

• QUERY

• VAR <name>

• STEM <name>

‘Results:’

• .SELECTED The following error codes can be returned in RC2:

• RXERR NOCURRENT

• RXERR ENVCHANGED

• RXERR NONETWORKSELECT

7.5.4.34 The SORT Command

‘Syntax:’ SORT [SORT1 <string>] [SORT2 <string>] [SORT3 <string>] [SORT4 <string>]

‘Argument template:’

SORT1/K, SORT2/K, SORT3/K, SORT4/K DFA — the address utility for the AMIGA

‘Function:’

The addresses are sorted according to the sort criteria given in the DFA-Preferences. Alternatively you may explicitely give the search criteria by which you would like your addresses be sorted.

‘Argument description:’

• SORT1 <string>

• SORT2 <string>

• SORT3 <string>

• SORT4 <string> The following keywords are supported for the Sort parameters: − SALUTATION − FIRST − NAME − CO − STREET − ZIP − CITY − STATE − COUNTRY − BIRTHDAY − PHONE − FAX − EMAIL1 − EMAIL2 − EMAIL3 − COMMENT

‘Results:’

The following error codes can be returned in RC2:

• RXERR ENVCHANGED

• RXERR NONETWORKSORT

7.5.5 Error Codes

Besides the usual error code that can be found in the RC variable^{2} the Arexx interface of DFA returns a second return value in the variable RC2, which describes in further detail what has gone wrong.

Bolow you’ll find the error numbers and their meaning: 2 cf. Arexx User’s Reference Manual, William S. Hawes, P.O. Box 308, Maynard, MA 01754

DFA — the address utility for the AMIGA

Symbolic Name Meaning

10  RXERR_NONETWORKAPPEND....Could not append file,

as the network mode is activated

11  RXERR_MODIFIED...........The address file has been changed
12  RXERR_NOFILENAME.........No filename has been supplied
13  RXERR_NOCURRENT..........No current address available
14  RXERR_NOFIRST............No first address available

15 RXERR_NOICONIFY..........Iconifying the Editor went wrong 16 RXERR_NONETWORKKILL......Deleting addresses is not allowed

while in network mode

17  RXERR_NOLAST.............There is no last address

18 RXERR_NOLOAD.............The address file could not be loaded

19  RXERR_NONETWORKNEW.......It is not possible to enter new

addresses in network mode

20  RXERR_NONEXT.............There is no next address

21 RXERR_NONEXTSEL..........There is no next selected address

22  RXERR_NOPREV.............There is no previous address

23 RXERR_NOPREVSEL..........There is no previous selected address 24 RXERR_SYNTAX.............The Arexx command contained an syntax error

25  RXERR_WRITEALL...........The "printing" of addresses (all)

into a file went wrong

26 RXERR_WRITESEL...........The "printing" of addresses (selected)

into a file went wrong

27 RXERR_WRITECURRENT.......The "printing" of the current address

into a file went wrong

28 RXERR_NOSAVE.............The address file could not be saved 29 RXERR_NONETWORKSAVE......Saving your address database is not

allowed in network mode

30 RXERR_NOTFOUND...........The wanted address could not be found 31 RXERR_NONETWORKCLEAR.....Deselecting addresses is not allowed

while being in network mode

32 RXERR_NONETWORKSELECT....Selecting addresses is not allowed

in network mode

33 RXERR_NONETWORKSORT......Sorting addresses is not allowed

while being in network mode

34 RXERR_NONETWORKEDIT......Editing addressen is not allowed

while being in network mode

35  RXERR_OPNPRT.............Could not open printer device
36  RXERR_CRTEXTIO...........Could not reach the printer

(ExtIO could not be initialized)

37  RXERR_CRTPORT............Could not reach the printer

(Could not open message port)

38  RXERR_NOMEM..............No more Ram available

39 RXERR_NORELOAD...........Reloading the address file went wrong

40  RXERR_NOWRITELOCK........could not get write access to the

address file

41 RXERR_NODIR..............Could not locate the needed drawer

42  RXERR_NOFILE.............Could not find the address file

43 RXERR_APPENDFAILED.......Appending an address file went wrong 44 RXERR_ENVCHANGED.........Execution of the current Arexx command

has been canceled, as the environment changed (e.g. new address file)

45 RXERR_NOEXTERNAL.........Trying to access the external file had no success

46  RXERR_NOSERIALDEV........Could not open serial device

47 RXERR_NOPHONENUM.........Current address doesn’t have a phone number

48  RXERR_SERNOCARRIER.......Phone: No Carrier
49  RXERR_SERNODIALTONE......Phone: No dialtone
50  RXERR_SERRING............Phone: Ring
51  RXERR_SERBUSY............Phone: Line is busy
52  RXERR_SEROK..............Phone: Modem sent ’Ok’

53 RXERR_SERERROR...........Phone: Modem recognized an error

54  RXERR_SERCONNECT.........Phone: Data connect
55  RXERR_SERVOICE...........Phone: Voice connect

56 RXERR_SERUNKNOWN.........Error caused by the serial device

Chapter 8: The DFA-Editor program Page 41

8 The DFA-Editor program

The DFA-Editor is the graphical user interface of DFA.

It can be started seperately or may be invoked (by the hotkey for example) out of the DFA-Server.

8.1 Tooltypes

If you click once on the DFA-Editor icon and then activate the Information program of the Workbench, you may adjust the following tooltypes:

• PUBSCREEN

If you enter the name of a public screen together with this tooltype, you may select an alternate (public) screen for the DFA-Editor to open on.

For example:

         PUBSCREEN=TERM

If the given public screen exists when the DFA-Editor comes up, it opens its window(s) on this screen, otherwise on the default public screen, usually the Workbench screen.

• UPPERPUBSCREEN

If you enter ‘UPPERPUBSCREEN=YES’, you tell the DFA-Editor to open its window(s) on the frontmost public screen. If the current frontmost screen isn’t public, the DFA-Editor comes up on the screen you selected with the tooltype PUBSCREEN; if this fails as well, the DFA-Editor opens its window(s) on the default public screen.

• GUIDEFILE

To make the online help work, the DFA-Editor has to know where to find the online help file. This can be done via the tooltype GUIDEFILE. This tooltype is usually already be set by the install program.

• DFAPREFS

As the DFA-Preferences program can be started directly out of the DFA-Editor, the DFA-Editor has to know where to find the DFA-Preferences program. The complete path name must be given with the tooltype DFAPREFS and is already set by the install program usually.

8.2 CLI parameter

If you start the DFA-Editor from the shell (or the CLI), it uses the tooltypes of the corresponding ‘.info’ file! However, you may overwrite these setting by giving one or more of the following CLI parameters:

     PUBSCREEN/K,UPPERPUBSCREEN/K,GUIDEFILE/K,DFAPREFS/K

This means, you may use…

… ‘DFAEditor PUBSCREEN=TERM’ to make the DFA-Editor open its window(s) on the public

screen named ‘TERM’. … ‘DFAEditor UPPERPUBSCREEN=YES’ to make the DFA-Editor open its window(s) on the front-

most screen, as long as this is a public screen. … ‘DFAEditor GUIDEFILE=Help:DFAEditor.guide’ to set another path for the online help file. … ‘DFAEditor DFAPREFS=C:DFAPrefs’ to set another path for the call of the DFA-Preferences

program.

8.3 The Main Window

The main window of the DFA-Editor can be resized using the sizing gadget on the lower right corner of the window. The window contents automatically adapts on the new window dimensions.

As the main window is a socalled application window you may drag the icon of an address file on this window and drop it, to let the DFA-Editor load the corresponding address file.

The main window of the DFA-Editor gives you the following possibilities:

• The address listview The address listview displays the stored addresses that are available as soon as you entered some (more details about this subject later on). This listview can be controlled by mouse as usual or completely by keyboard:

Cursor down The next address becomes the current one.

SHIFT+Cursor down The list scrolls up by one page.

CRTL+Cursor down The last address becomes the current one.

Cursor up The previous address becomes the current one.

SHIFT+Cursor up The list scrolls down by one page.

CRTL+Cursor up The first address becomes the current one.

The current entry is selected. SPACE To make this state visible, the current entry is leaded by a ‘>’ character.

RETURN Activates the Full window, as long as there is an active address item.

SHIFT+RETURN Activates the Edit wenster, as long as there is an active address item.

SHIFT+SPACE Set the start of the block or the end of the block respectively

You may select entries of the address list with Shift Left Mousebutton or Middle Mousebutton Left Mousebutton (3-Button mouse needed); alternatively you may hit SPACE on the keyboard. If you hit RETURN, it has the same effect as if you had selected the Full button: The current address is displayed in another window in further detail (please see Section 8.6 [Full], page 51 for details). Synonymously you may hit SHIFT RETURN to envoce the Edit window (see Section 8.7 [Edit], page 53). SHIFT RETURN sets the start of the block or — if you’ve already set it previously — the end of the block. After the block end is set, the selection state of all items between start end end of the block is toggled, i.e. the selected addresses become deselected and vice versa. The same result can be reached, if you use the left mouse button together with the SHIFT qualifier. Please notice that you may change the fields of the addresses that are displayed in the address listview using the DFA-Preferences program.

• Der Panel–Listview Below the address listview you find another listview. It is used to display some more fields of the current address item. If you use the default settings you’ll see the phone number, the three email addresses and the comment. However, by using the DFA-Preferences program you may change the contents of these lines! As this listview doesn’t have a title and for this reason no shortcut either, you may scroll it using the following keys:

ALT+Cursor down The list scrolls up by one line.

ALT+Cursor up The list scrolls down by one line.

Of course, the list can only be scrolled if there is 1st an active entry and 2nd more lines to display than visible at the same time. The contents of this listview can be configured in quite a wide range (see DFA-Preferences: Panel Listview, for details).

Just below the the address listview there are four buttons that may be used to change the selection of the displayed addresses:

• All Selects all displayed addresses.

• By name A window is opened in which you may enter a search pattern (case insensitive!). After you hit Ok (or RETURN) all entries of the address list are selected that match the pattern. The difference to the Search function is that 1st you can’t toggle case sensitivity and 2nd all address fields are searched through. Furthermore all matching items are marked and the first matching item is activated. A detailed description of the allowed search patterns can be found in Section 8.9 [Search], page 54.

• Clear All selected addresses are deselected.

• Jump to The next selected entry becomes the current one, as long as there is at least one selected entry left.

At the lower end of the main window of the DFA-Editor you can find the gadgets Quit and

Remove.

If you hit Quit the DFA-Editor is terminated! The DFA-Server, if it is currently running, is not ended however, so you can still reach the DFA-Editor by a doubleclick on the application icon and/or the hotkey (default: Left Amiga F5)

If you select Remove, not only the DFA-Editor is ended, but although a currently running DFA-Server.

If you terminate the DFA-Editor using Quit or Remove, the addresses are saved if they had been modified before. This behaviour can be customized in the DFA-Preferences however.

8.4 Tool Gadgets

It is possible reach the socalled tool gadgets from the DFA-Editor. Using these gadgets you can easily reach the following functions :

• Full Shows the current entry in another window in great detail; see Section 8.6 [Full], page 51, for details.

• Edit If you hit Edit, you may change the current entry. If there is no current address, this button is displayed "ghosted" and can’t be selected (cf. Section 8.7 [Edit], page 53).

• New If you select New, you may enter a new address. The functions within this window are exactly the same as within the Edit window; see Section 8.7 [Edit], page 53 for details.

• Kill Kill can either be used to delete the current entry (if there is one) or all marked entries. To make sure that nobody deletes precious addresses, a safety requester pops up!

• Search Another window pops up, in which you may enter a search string as well as some other search related settings (cf. Section 8.9 [Search], page 54).

• Search Next Search Next can be selected only, if there has been given a search pattern previously using the Search function. If this is the case, the next matching entry is displayed (not marked!).

• Print After you have selected Print, another window pops up, which gives you several ways to print your addresses, as it is described in Section 8.10 [Print], page 56.

• Sort Sort sorts the complete address list. As long as you use the default setting the list is sorted by name, first name and city (in this order). If you prefer another sort order, you may change it in the DFA-Preferences.

• Dial Another window pops up, which allows you to dial phone numbers (cf. Section 8.11 [Dial], page 57).

Please notice that there is a corresponding menu item for every tool gadget. All functions can therefore be reached by keyboard shortcuts as well!

8.5 The Menu Items

8.5.1 About

The About window gives you some useful informations about the program and the author:

• The version number and a corresponding remark, if you got a registered version.

• The copyright notice and the author’s address (incl. email addresses).

• Your name and your address, as well as the serial number of the program, if you are a registered user.

8.5.2 Info

The window which pops up as soon as you select the Info menu item gives you some useful informations:

• Arexx Port In this field you can look up the name of DFA’s Arexx port. Usually this is ‘DFA’. However, you can change the portname by changing the PORTNAME tooltype (of the DFA server program!) If you have already started a program that uses the portname DFA, DFA adapts the portname by appending "numbers" (e.g. ‘DFA.1’, ‘DFA.2’,…) until a clear portname is found.

• Addresses The number of the currently loaded address is displayed here. The first number (‘d’ means displayed) is the number of actually displayed entries, the second number (‘a’ means available) is the number of addresses available.

• Memory Condition The three fields give a quite detailed report about the current memory condition.

8.5.3 New file

If you want to create a new, i.e. empty address file, you may use this menu item. A new address file named unnamed.dfa is created then. You may use the Save as... menu item to save this address file under another name.

Be careful:

Please do never use Save as default after you created a new file to save it, as the new address file (the empty one!) would overwrite your old "default" address file, which means that the old addresses would have been lost!

8.5.4 Open

Use Open to load another address file. The currently loaded addresses are not kept in memory but replaced by the loaded! After you selected Load, a file requester pops up, which lets you enter the address file. Notice that only files can be loaded, which have the proper format — files that have been created by DFA before! Files with the wrong format can’t be loaded by DFA and the already loaded addresses keep unchanged!

8.5.5 Append

Append allows you to append further addresses to the existing ones. The already loaded addresses are not deleted before. Be careful! If you append two identical address lists, every entry will appear twice, which is usually not desired.

It is possible to set the group selection for the addresses that are to be appended in the little window that appears after you selected Append. These group selection flags are set to all addresses that shall be loaded!

Please notice that already existing group flags of addresses in the file to append are not erased! This means that you may only add new group flags to the addresses but not unset existing ones!

Ok accepts the set group selection and appends the selected address file; Cancel not only cancels the group selection but the complete appending procedure, i.e. no addresses are appended!

8.5.6 Save

Save saves the addresses under the current name. The current name is either the one you entered in the preferences (default: ‘s:adr.file’) or the one you selected before when loading an address file.

8.5.7 Save as…

Save as saves the address as well. However, you are able to select a filename before (a file requester pops up). This may be useful, if you have several address files for different purposes or you’ve created a new address file just before (cf. Section 8.5.3 [New file], page 46) and want to enter a new name for your address file.

8.5.8 Save as default

The addresses will be saved under the name specified in the preferences (default: ‘s:adr.file’), if you select Save as default, even if you loaded a file with another name before.

Be careful:

If you have loaded an address file with another but the default name before (or you’ve created a new address file), you should not use Save as default in any case, as the current address file would overwrite the default address file.

8.5.9 Quit

Quit ends the DFA-Editor! The DFA-Server, if it is currently running, is not ended, so it the DFA-Editor still can be reached by a double click on the application icon or by the hotkey

(Default: Left Amiga F5).

8.5.10 Remove

Select Remove to end not only the DFA-Editor but a currently running DFA-Server as well.

If you end the DFA-Editor using Quit or Remove DFA, the addresses are saved, if they have been modified before. This behaviour can be customized in the DFA-Preferences program however.

8.5.11 Full

Shows the current entry in a separate window. All fields of the address are displayed here; See Section 8.6 [Full], page 51.

8.5.12 Edit

Using Edit gives you the chance to change the current address item, if there is one. If there is no current entry, the corresponding button is "ghosted" and can’t be selected (see Section 8.7 [Edit], page 53).

8.5.13 New

Select New to enter a new address. The functions inside of this window are exactly the same as within the Edit window; see Section 8.7 [Edit], page 53 for details.

8.5.14 Kill

Use Kill to delete either the current entry (if there is one) or all marked addresses. To make sure that nobody deletes precious addresses, a savety requester pops up.

8.5.15 Copy to Clipboard

This function makes it possible to copy the current address to the clipboard (unit 0). Every other program that supplies clipboard support, can access these data.

8.5.16 (Un)set Groups

After you have selected this menu item, another window pops up that allows you to change the group flags of one, several or all addresses fast and comfortably. Using the upper eight group checkboxes, you can select, which group flags you want to have cleared. The lower eight group checkboxes can be used to set the corresponding group flags. To control even these gadgets by keyboard, you have to push the CRTL key in addition to the shortcut itself. Please notice that these actions take effect only to the active addresses, i.e. the addresses that can be seen within the main address listview. Hit Active to change the current address (if there exists one), Marked to change all marked addresses and All to (un)set the group flags of all (visible) addresses.

8.5.17 Search

Another window pops up that allows you to enter several search criteria (cf. Section 8.9 [Search], page 54):

8.5.18 Search Next

Search Next can be selected only, if there has been given a search pattern previously using the Search function. If this is the case, the next matching entry is displayed (not marked!).

DFA — the address utility for the AMIGA

8.5.19 Print

After you have selected Print, another window pops up, which gives you several ways to print your addresses, as it is described in Section 8.10 [Print], page 56.

8.5.20 Sort

Sort sorts the complete address list. As long as you use the default setting, the list is sorted by name, first name and city (in this order). If you prefer another sort order, you may change it in the DFA-Preferences.

8.5.21 Dial

Another window pops up, which allows you to dial phone numbers (cf. Section 8.11 [Dial], page 57).

8.5.22 Preferences

The DFA-Preferences program is loaded. Please notice that the tooltype DFAPREFS of the DFA-Editor has to be set correctly to enable the DFA-Editor to find the DFA-Preferences program! This has usually already been done by the install program.

8.5.23 Edit Template

The window that pops up will be quite familiar to you, as it is exactly the same as in Section 8.7 [Edit], page 53 or Section 8.8 [New], page 54. You may enter several presets here, which are automatically copied into the corresponding fields, just before you enter a new address. The template strings are saved in the address file, so it is possible to have different templates for every address file.

8.5.24 Write Access

This menu item can only be selected, if you have activated the network mode before.

Starting with version 2.0, DFA supplies a simple network mode. This network mode has been developed to make it possible that several connected computers can use the same address file.

The problem that comes up, when an address file is used by more than one user at the same time is the following:

If several users change the address file at the same time, the last saved address file overwrites the previously saved ones. The previously saved changes are lost!

DFA tries to solve this problem in the following way:

If the network mode is active, all operations that change the address database are forbidden, i.e. saving, changing, marking of addresses is not possible. It is allowed however, to view the addresses (‘Full’), to scroll within the address list, to print the addresses and so on.

If you want to do some changes on the addresses, you have to obtain write access to the address file. This can be done using the Arexx command ‘ATTEMPTLOCK’ or by selecting the menu item ‘Write Access’ in the DFA-Editor.

If your try to get write access was successful, you may change the addresses. If you’ve completed your changes, you should free the write access using either the Arexx command ‘FREELOCK’ or the menu item ‘Write Access’ in the DFA-Editor. After you have done so, other members of the Lan can get write access to this address file.

Please note that you will only get advantages out of the network mode, if you have really installed a network and access the same address file from different computers! Activating the network mode on a computer that is not connected to a network has no advantages at all!

8.5.25 Save Preferences

As the window coordinates and several other settings of the DFA-Editor are saved together with the preferences that can be tuned in the DFA-Preferences program, there is the opportunity to save the preferences not only from within the DFA-Preferences program, but from the DFA-Editor as well using this menu item.

8.6 Full

Full can only be selected, if there is a current address! If this is the case, another window pops up (title: Full address…), which displays all fields of the current address.

If there is an external file for the current address, you may select the View external button to start the ascii viewer together with the corresponding external file.

Please notice that the ascii viewer of your choice may be entered in the DFA-Preferences program (cf. Section 8.7 [Edit], page 53).

The listview that displays the different fields of the current address may be scrolled as follows:

Cursor down The listview scrolls up.

Cursor up The listview scrolls down.

Below this listview there is a gadget panel that may be used to step through the addresses:

• First The first entry of the address list is displayed. Besides the underlined letter you may alternatively use CRTL + Cursor Left as the corresponding gadget shortcut.

• Last The last entry of the address list is displayed. Besides the underlined letter you may alternatively use CRTL + Cursor Right as the corresponding gadget shortcut.

• Prev. Sel. If there is a selected item infront of the current one, it is is displayed. Besides the underlined letter you may alternatively use SHIFT + Cursor Left as the corresponding gadget shortcut.

• Next Sel. If there is a selected item behind of the current one, it is is displayed. Besides the underlined letter you may alternatively use SHIFT + Cursor Right as the corresponding gadget shortcut.

• Prev.

The previous entry of the address list is displayed. Besides the underlined letter you may alternatively use Cursor Left as the corresponding gadget shortcut.

• Next The next entry of the address list is displayed. Besides the underlined letter you may alternatively use Cursor Left as the corresponding gadget shortcut.

Ok closes the window, Edit opens the Edit window. The Edit window is described in further detail in Section 8.7 [Edit], page 53.

DFA — the address utility for the AMIGA

8.7 Edit

If there exists a current address (cf. Section 8.6 [Full], page 51), a window pops up, where you may change any part of the address. The string gadgets have some additional features compared to the "original" ones:

− Hit RETURN to end the input to the current gadget and go to the next string or integer gadget. You may use the ENTER key alternatively.

− Hit SHIFT RETURN to end the input to the current gadget and go to the previous gadget.

− CURSOR UP or CURSOR DOWN brings the cursor to the previous or next string gadget respectively.

− Hit ESCAPE to cancel the input completely. The window which contains the gadget is closed and all previous input is discarded.

− If the cursor stands within a string gadget, the gadget shortcuts can be reached by hitting RIGHT AMIGA–letter and SHIFT. If you don’t hit SHIFT additionally, the keyboard shortcut is used either by the internal string gadget edit features (like RIGHT AMIGA X or RIGHT AMIGA Q) or by additional tools (like NewList).

There exist the following fields:

Address  Address

First First name

Name  Name

c/o may be used for additional information Street Street/No.

Zip  Zip
City  City
State  State
Country  Country
Birthday  Birthday

Phone Telphone number Fax Telefax number Email1 email address # 1 Email2 email address # 2 Email3 email address # 3

Comment  Comment

Please notice that it is possible to enter more than one phone number into the Phone field. These numbers can be dialed correctly as well, as long as you separate them by a pipe sign (|).

Using the Address selected checkbox you may set (or unset) the selection state of the current address.

If you hit External you activate the ascii–Editor (default: ‘ed’). This makes sense if you want to enter a long text (for example: How to find the way to Joe User, living in New York City).

When you have finished your inputs, save the file with the given name and quit the editor. The DFA-Editor remembers the file name, so it is possible from now on, to access this external file e.g. from the Full window (cf. Section 8.6 [Full], page 51) or from the Edit window to modify it.

As soon as the there is an external file for the current address item, the Del. button can be reached that makes it possible to delete an external file as soon as you no longer need it.

All external files get their unique file name by DFA. This file name is guarantied to be unique within the given directory (default: ‘s:DFAExternals’). DFA only saves the filename — not the file itself, so this is completely up to you. Please do not change the filename given by DFA, as if you do so, DFA will no longer be able to find this file! If you want to delete one (or more) external files, you have to use the Del. function of this window. This is the only way to make sure that not only the file itself is deleted, but the filename, which is stored by DFA as well!

You may set your favourite ascii editor as well as the wanted directory for the external files in the DFA-Preferences program.

Hit Ok to use the changes; if you select Cancel instead, all changes are discarded.

8.8 New

New opens a window that looks exactly like the Edit window! The templates (cf. Section 8.5.23 [Edit Template], page 50) have been copied into the corresponding fields and you may enter a new address now. All functions within this window are the same as in the Edit window (see Section 8.7 [Edit], page 53 for details).

8.9 Search

You may enter several search criteria in this window.

• Search for Here you may enter the search pattern. You may (and must) use the known AMIGA patterns. It is needed to enter infront of and behind the search string either a ‘#?’ or a ‘*’ if you want to look for partial strings! This means, you have to input ‘#?PartialString#?’ instead of ‘PartialString’. Let’s say you want to search for first names that start with ‘Mi’, you have to enter ‘Mi#?’. If you want to look for cities that end in ‘town’, you have to enter ‘#?town’. The following patterns are supported:

? The question mark matches any single character. For example, the pattern matching string ‘A?B’ matches any string that is three letters long, that starts with an ‘A’ and ends with a ‘B’.

# The number sign matches strings containing zero or more repetitions of the expression that immediately follows the ‘#’ in the pattern matching string. For example, the pattern matching string ‘#A’ matches any string that consists of one or more of the ‘A’ character. The pattern matching string ‘#?’ matches any non-NULL string. The ‘#’ can apply to entire substrings delimited by parentheses. For example, the pattern string ‘#(AB)’ matches any string consisting of one or more repetitions of the substring ‘AB’ (‘AB’, ‘ABAB’, ‘ABABAB’,…).

% Matches the NULL string.

| This is the OR symbol. This matches strings that contain the expressions on either side of the OR sign. The expressions and the OR symbol need to be enclosed in parentheses. For example, the pattern matching string ‘(A|B)’ matches the string ‘A’ or the string ‘B’. The pattern matching string ‘A(B|%|C)’ matches the strings ‘AB’, ‘A’, and ‘AC’.

~ The tilde negates the expression that follows it. All strings that do not match the expression that follows the tilde will match the expression with the tilde. For example, the pattern matching string ‘~(#?.info)’ matches any string that does not match the string ‘#?.info’ (does not end with the substring ‘.info’).

* The star is provided as an synonym to ‘#?’.

[ ] All characters within brackets indicate a character class. Any character in the character class qualifies. Within a character class, a character range can be indicated by specifing the start and stop character, separated with a minus sign. Note that character classes are case sensitive. If character classes are to be used in a case insensitive form, they should be translated to uppercase. Here are some examples:

[ACF]#? matches strings starting with ‘A’, ‘C’, or ‘F’

[A–D]#? matches strings starting with ‘A’, ‘B’, ‘C’ or ‘D’

[~ACF]#? matches strings not starting with ‘A’, ‘C’ or ‘F’

’ The quote character neutralizes the special meaning of a special character. Here are some examples:

’#’? matches only the literal string ‘#?’ ’?(A|B|C|%)’# matches the literal strings ‘?#’, ‘?A#’,

          ‘?B#’, ‘?C#’

” The quote itself

( ) Parentheses group special characters. The expression within the parentheses is a subpattern.

• Ignore case If the box is "checked", the search will not be case sensitive. If the box is empty however, case will be noticed.

• All Select all fields.

• None Deselects all fields.

• Fields to search You may select which fields you would like to search through. It is possible to toggle each single field. The default is to search trough all fields.

• Search Search starts the searching. If a matching item is found, the address listview scrolls to the corresponding place and the matching entry becomes the current one.

• Cancel Cancel terminates the search.

8.10 Print

This window contains all functions that may be used to print the addresses.

• Print You can choose between all (All), the marked (Selected) and the active address (Active). If there is an active item, it will be shown in the Active Item field.

• Print to You can either print to the printer (PRT:) or to a file (File). If you print your addresses to a file, the settings below don’t take effect. Notice that always all parts of an address are written to the file!

• Print style You may choose, whether the addresses are to be printed in NLQ or Condensed style (or both). If you select Formfeed, a formfeed is performed after the very last printed address entry. To switch this off will be useful, if you want to print a few address labels only and you don’t want to waste these sheets.

• Print form There are five different forms, between which you may choose: − Short address list − Full address list − Phonelist (fixed columns)

DFA — the address utility for the AMIGA − Mailing labels (70mm x 36mm) − Address cards

• Print If you did select File for output, a file requester will be opened, which allows you to enter the filename. Be careful! This output file is not compatible to the internal address file format (default: ‘s:adr.file’); so don’t overwrite this (these) file(s). To write the addresses to a file may be useful if you want to export addresses, but don’t wan’t to use Arexx. If you send your addresses to the printer (PRT:), another window is opened that displays the progress of printing. You may stop the output after every complete address (Cancel). If only the active item is printed, no progress indicator is shown.

• Cancel Cancel terminates the complete printing process.

Please notice:

Of course the possibilities of DFA to print addresses are a bit limited and don’t have any chance to compete with the facilities given by e.g. a dtp program. I think however that it does make much more sense to let programs print your addresses that are specialized in printing. Such programs give you much more features and often result in a much better printing quality than an address utility could ever reach. So DFA should be used to supply the stored addresses to (spezialized) printing programs and not not print the addresses by itself. An ideal mean for this purpose is the Arexx port for sure, which is not only part of DFA, but of many printing programs as well.

Even if your favourite printing program doesn’t have an Arexx port, you may use DFA’s Arexx interface to export and format the address, so other programs can easily use them.

8.11 Dial

For owners of modems only!

If you have a modem connected to your computer, it will be possible to let the computer dial the phone number(s). If you have entered more than one number, it is strictly recommended that you seperate them by a pipe sign (|). When you start dialing (hit Dial), a window pops up that displays who is called. You can cancel the dialing at every time (Cancel) or skip the current number (Skip). The modem settings can be changed in the DFA-Preferences program.

9 The DFA-Preferences Program

The DFA-Preferences program (filename: ‘DFAPrefs’) may be used to configure DFA. It can be started from the DFA-Editor (menu item ‘Settings->Preferences’) as well as directly from the Workbench or the Shell.

9.1 Tooltypes

If you click on the DFA-Editor icon once and then activate the Information program of the Workbench, you may adjust the following tooltypes:

• PUBSCREEN

If you enter the name of a public screen together with this tooltype, you may select an alternate (public) screen for the DFA-Preferences program to open on. For example:

         PUBSCREEN=TERM

If the given public screen exists when the DFA-Preferences program comes up, it opens its window(s) on this screen, otherwise on the default public screen, usually the Workbench screen.

• UPPERPUBSCREEN

If you enter ‘UPPERPUBSCREEN=YES’, you tell the DFA-Prefs program to open its window(s) on the frontmost public screen. If the current frontmost screen isn’t public, the DFA-Preferences program comes up on the screen you selected with the tooltype PUBSCREEN; if this fails as well, the DFA-Preferences program opens its window(s) on the default public screen.

• GUIDEFILE

To make the online help work, the DFA-Preferences program has to know where to find the online help file. This can be done via the tooltype GUIDEFILE. This tooltype is usually already set by the install program.

9.2 CLI–Parameter

If you start the DFA-Preferences program from the shell (or the CLI), it uses the tooltypes of the corresponding ‘.info’ file! However, you may overwrite these setting by giving one or more of the following CLI parameters:

Chapter 9: The DFA-Preferences Program Page 59

 PUBSCREEN/K, UPPERPUBSCREEN/K, GUIDEFILE/K

This means, you may use…

… ‘DFAPrefs PUBSCREEN=TERM’ to make the DFA-Preferences program open its window(s) on

the public screen named ‘TERM’. … ‘DFAPrefs UPPERPUBSCREEN=YES’ to make the DFA-Preferences program open its win-

dow(s) on the frontmost screen, as long as this is a public screen. … ‘DFAPrefs GUIDEFILE=Help:DFAPrefs.guide’ to set another path for the online help file.

9.3 The DFA-Preferences Main Window

From the main window you can reach all subwindows. You may either open one subwindow at a time or several subwindows at the same time.

Furthermore the following functions may be reached via buttons:

• Save Saves the preferences.

• Use Saves the preferences temporarily to ‘ENV:’, i.e. the preferences stay active until you reboot the computer.

• Remove DFA Removes a running DFA application, i.e. the DFA-Server and the DFA-Editor will be terminated, if they are currently running.

• Quit Quits the DFA-Preferences program.

You may reach the following functions via the Project menu:

• About In the About window you’ll find something useful about the program itself and the author: − the version number as well as a notice, if you are a registered user. − The copyright notice and the address of the author (incl. email address(es)). − Your name, address and the serial number if you are a registered user.

• Save See above.

DFA — the address utility for the AMIGA • Remove DFA

See above.

• Quit See above.

9.4 Miscelleneous

9.4.1 Application Icon

The string gadgets XPos and YPos make it possible to enter the application icon’s position on the Workbench screen (referring to the upper left corner of the screen) Alternatively to this direct input of raw coordinates there is a more convenient possibility to change the application icon’s position: As soon as you click (once) on the cycle gadget Open position window, a small window (title: Move me) is opened. This window can be moved (using the mouse) on the Workbench. As soon as the position fits your wishes you should click on the gadget Close position window (the same gadget as before, just with another text in it). The Move Me window is closed and the new icon position is displayed in the XPos and YPos gadgets. Using the checkbox gadget Show, you can switch the application icon on and off. Furthermore it can be used to make a previously changed icon position visible, by switching the application icon off and then on again.

9.4.2 Auto Save

There are several ways to make sure that your addresses are saved regularly:

• Save always If you have changed some addresses using the DFA-Editor, the address file is saved as soon as you leave the DFA-Editor — no matter if using Quit or Remove.

• Ask If the addresses have been changed, a requester pops up as soon as you want to leave the DFA-Editor, which asks you if you would like to save the addresses.

• Remove only The changed addresses will only be saved, if the DFA-Editor is left using the Remove function (i.e. you quit the complete DFA application), but not if you use the Quit function.

DFA — the address utility for the AMIGA Be careful: Please notice that the functions Quit and Remove of the DFA-Editor are exactly the same, if you don’t have the DFA-Server program running at the same time! If there is running only the DFA-Editor and you’ve choosen Remove only for auto save mode, changed addresses are not saved, if you select Quit in the DFA-Editor.

9.4.3 Icon Creation

If you want DFA to save an icon together with the address files, you have to switch this checkbox to ON state. To get an icon for every address database file make sense e.g. if you like to use DFA’s application window.

9.4.4 Secret Mode

Not everyone likes DFA informing everybody about how many addresses are sorted, saved, loaded and so on. If you select Secret Mode, instead of the number of addresses, three dots are displayed.

9.4.5 Networking

Starting with version 2.0, DFA supplies a simple network mode. This network mode has been developed to make it possible that several connected computers can use the same address file. The problem that comes up, when an address file is used by more than one user at the same time is the following:

If several users change the address file at the same time, the last saved address file overwrites the previously saved ones. The previously saved changes are lost!

DFA tries to solve this problem in the following way:

If the network mode is active, all operations that change the address database are forbidden, i.e. saving, changing, marking of addresses is not possible. It is allowed however, to view the addresses (‘Full’), to scroll within the address list, to print the addresses and so on.

If you want to do some changes on the addresses, you have to obtain write access to the address file. This can be done using the Arexx command ‘ATTEMPTLOCK’ or by selecting the menu item ‘Write Access’ in the DFA-Editor.

If your try to get write access was successful, you may change the addresses. If you’ve completed your changes, you should free the write access using either the Arexx command ‘FREELOCK’ or the menu item ‘Write Access’ in the DFA-Editor. After you have done so, other members of the Lan can get write access to this address file.

• Network mode The checkbox gadget Network mode may be used to switch the network mode on and off. Please notice however that using this feature makes only sense, if you really have a network installed and want to access the same address file from several computers.

• Retries Number of retries to get write access to the current address file.

• Retry delay Number of seconds that may pass between two tries to get write access.

9.5 Address Listview

This window may be used to configure the main (address) listview of the DFA-Editor:

Each line of the address listview contains one address. Each line may have up to eight columns. Each column can display any field of the particular address.

• Column This cycle gadget supplies a mean to select the desired column. You may configure up to eight columns, i.e. ‘Column 1’ through ‘Column 8’.

• Fields This listview contains the field name of all address fields that may be used. Please notice that the ‘Groups’ entry displays all eight group flags of the particular address using the following format:

     12345678

…if there are all groups flags set. For every group that isn’t set, instead of the corresponding group number there will be displayed a minus sign (‘-’). Some examples:

123----- Only the groups 1,2 and 3 are set 1-3-5--- Only the groups 1,3 and 5 are set ------78 Only the groups 7 and 8 are set -------- No group flag is set

• Weight The number given here is the percentage of the particular column referring to the complete line. Please notice that it is possible to give up to 100 %, but not more than this! If there

DFA — the address utility for the AMIGA is not enough space left for the column you’re currently working on, you have to reduce one or more columns accordingly. If you enter a number into the Weight gadget that is too large, the program corrects the value and replaces it with the highest possible number. As the display of the group flags needs a fixed number of eight characters, it is not possible to give a weight for this field.

• Free and Used To give you a general view of the used and free space within the line, there are two gadgets, Used [%] and Free [%], which display the used and free space respectively.

• Clear Column To clear a column it is not sufficient to set its weight to ‘0’! If you want to clear a column, you have to select the desired column (using the Column gadget) and then hit the Clear column gadget. Please notice that "empty" columns are not possible! Only the columns in front of the first cleared column are displayed. Example:

You have set the columns 1,2,3 and 4 to the fields ‘Name’, ‘FirstName’, ‘Phone’ and ‘Town’. If you clear column 3 (‘Phone’) and select Ok, only columns 1 and 2 (‘Name’ und ‘FirstName’) are still visible.

• Ok und Cancel With Ok you accept the changes you’ve done; if the DFA-Editor is currently running, the changes are applied to it and are visible at once. Cancel discards the changes.

This all may have sounded a bit complicated, so I’ll give you an example that shall make clear in a step by step procedure, how to configure your "own" format for the address listview:

Let’s say, you want the following format of the listview’s lines:

      <First Name> <Name> <Phone> <Group Flags>
          30%        30%    40%        ---

…then you should do like this:

Click on the Columns gadget until ‘Column 1’ is shown. Now select ‘First name’ in the Fields listview and enter ‘30’ into the Weight string gadget. Click once on the Columns gadget — ‘Column 2’ appears. Select ‘Name’ in the Fields listview and enter ‘30’ into the Weight string gadget. Click on the Columns gadget once again

— ‘Column 3’ appears — and select ‘Phone’ from the Fields listview. Please enter ‘40’ into the Weight string gadget. One more click on the Columns gadget brings ‘Column 4’ to front. Now select ‘Group flags’ from the Fields listview. Please notice that the Weight gadget is displayed "ghosted" now and therefore no input into this gadget is possible; this is not needed either as the group flags always need eight characters. Please click on the Columns gadget one more time to make ‘Column 5’ visible. Select Clear column to make sure that the 5th column is "clear". Hit Ok to make the changes active — if the DFA-Editor program is currently running, all changes will be visible at once.

9.6 Panel Listview

Below the address listview of the DFA-Editor main window there is another listview, called "Panel listview". This is used to make further information about the current address visible. It is possible to display up to 16 lines. Every line can display any field of the address; the line title may be changed as well.

Furthermore it is possible to adjust the number of lines that are visible at a time.

• Lines Every click on the Lines cycle gadget makes the next line number visible (‘Line 1’ through ‘Line 16’). After the last line number (16) ‘Line 1’ is displayed again.

• Fields This listview corresponds to the one, you already got to know in Section 9.5 [Address Listview], page 62. Using this listview, you may select, which field of the particular address you want to get displayed in the currently selected line (cf. Lines).

• Title It is possible to give any line title you like; for example ‘Phone #:’ for the ‘Phone’ field or

‘Groups:’ for the ‘group selection’.

If you want e.g. that all three email address are displayed among one another (i.e. Line 1 through line 3), it could be a good idea, to give the first line (‘EMail 1’) the title ‘EMail:’ and leave line 2 and line 3 without a line title.

• Clear line Similarly to Section 9.5 [Address Listview], page 62 (see Clear Column) is is needed here as well to explicitly delete lines you don’t want to use any longer. Please notice that only lines infront of the first "cleared" line are displayed, even if there are lines with contents behind the first cleared line. This means that empty lines are not possible.

• Visible lines It is possible (within certain values) to change the size of the panel listview using this slider gadget. You may have up to 12 lines visible at a time! Please notice that the DFA-Editor reduces the number of visible lines automatically, if there is not enough space for the desired listview lines in the window.

• Ok and Cancel Using Ok you accept the changes and a currently running DFA-Editor program is updated at once. Cancel discards the changes.

The following example may help you to understand the Panel listview configuration more easily:

Let’s say you would like to get the following format in your Panel listview

              <Title>    <Contents>
     <Line 1> EMail:     [EMail 1]
     <Line 2>            [EMail 2]
     <Line 3> Comment:   [Comment]

Terms in ‘[’ and ‘]’ describe field names, terms in ‘<’ and ‘>’ serve as description.

First click as long on the Lines cycle gadget, until ‘Line 1’ appears (after the window has opened Line 1 is usually visible already). Now select ‘EMail 1’ from the Fields listview and enter ‘EMail:’ into the Title string gadget. Now click on the Lines gadget once — ‘Line 2’ should become visible. Select ‘EMail 2’ from the Fields listview and delete the contents of the Title string gadget, if it isn’t empty already. One more click on the Lines gadget brings ‘Line 3’ to front. Select ‘Comment’ from the Lines listview and enter ‘Comment’ into the Title string gadget. After one more click on the Lines gadget please click on the Clear line gadget, to delete line four if it isn’t empty. Hit Ok to accept the changes. If the DFA-Editor is currently running, the changes will be made visible at once.

9.7 Groups

First of all a few words about the group flags and their purpose and meaning:

It is possible to separate your addresses into groups and set the different group flags for every address item. Let’s say you’re using the first three (out of a maximum of eight) group flags and you’ve named them ‘Private’, ‘Business’ and ‘Other’ (see Section 9.7 [Groups], page 65, for more information).

Now you can set the ‘Private’ flag to all private addresses, the ‘Business’ flag to all business related addresses and the ‘Other’ flag to all address that don’t belong to any of these groups. Of course, you may set more than one flag at a time, if an address is a private one and business related as well, for example.

In the main window of the DFA-Editor you can select the groups you would like to see within the address listview (or use the Arexx command ‘CHANGEGROUPS’ accordingly). If you want to see the private and the business addresses, please activate both checkbox gadgets. Please notice that all actions, as for instance printing and searching is performed to the active (i.e. displayed) addresses only!

In the Groups window of the DFA–Preferences (cf. Section 9.7 [Groups], page 65) you may set which kind of relation between the groups you like. If you have activated more than one group at the same time (in the example above ‘Private’ and ‘Other’), you may select either the relation ‘via 'OR'’ or the relation ‘via 'AND'’:

• via ’OR’ Every address that has at least one of the wanted flags set, is displayed. If you activate ‘Private’ and ‘Other’, all addresses will be displayed that have either set the ‘Private’ or the ‘Other’ flag or both flags.

• via ’AND’ Only those addresses are shown that have set at least the groups you selected in the DFA-Editor main window. If the ‘Private’ and the ‘Other’ flags are both set, only addresses will be shown that have at least the ‘Private’ and the ‘Other’ flag set; addresses that have set only one of these flags are not shown.

Groups that don’t have any group flags set, can be made visible by deselecting all the group flags in the DFA-Editor main window (i.e. all check boxes are "empty").

Lets shift to the settings that are possible here:

• Groups

This cycle gadget allows you to set the group (‘Group 1’ through ‘Group 8’) you want to modify.

• Title

The name of the group can be entered into this string gadget. The desired keyboard shortcut has to be marked with a leading underscore (_), for example:

Group _1 The group’s name is ‘Group 1’, its shortcut is the digit 1

Pri_vate ‘Private’ is used for the group name, the keyboard shortcut is the v

Other (_3) ‘Other (3)’ is the group’s name, the digit 3 its keyboard shortcut

Please notice: Even if it is possible to use any keyboard shortcut you like, as long as it is an "usual" letter, you should remember that DFA does not check if the keyboard shortcut you enter is unique within the DFA application. Therefore you have to make sure by yourself that the keyboard shortcuts, you want to use, are not already be used by other gadgets. Please notice as well that the shortcut for the different groups are not only used within the DFA-Editor main window, but have to be unique in e.g. the DFA-Editor Edit window as well. The shortcuts, which are guarantied to be unique within the complete DFA application are the digits 1 through 8, which is the default as well.

The length of the group titles is limited to 15 digits, to make sure that no title overwrites the window contents and/or border.

• Group handling This cycle gadget allows you to select the kind of relation of several group flags (see above for details).

• Ok und Cancel After you’ve hit Ok, the changes become active, if you hit Cancel instead, all changes are lost.

9.8 Fonts

The fonts that shall be used by the DFA-Editor and the DFA-Preferences program can be set in this window. If you don’t change the default setting (Use System Font for every font), all program parts use the fonts you set in the Font Preferences program of the Workbench. Alternatively you may set any font you like for each of the four fonts used.

• DFA-Editor: Gadget font Here you may enter the font, which is used for the gadget texts of the DFA-Editor.

• DFA-Editor: List font For the different listviews of the DFA-Editor a seperate font may be chosen. Please notice that only non–proportional fonts are allowed here.

• DFA-Preferences: Gadget font see above, DFA-Editor: Gadget font.

• DFA-Preferences: List font see above, DFA-Editor: List font. Please notice, however that even proportional fonts are allowed here!

9.9 Modem

The settings that can be done here, have only a meaning for you, if you have a modem connected to your computer and plan to use DFA to dial the stored phone numbers.

Furthermore you should have a phone connected to your modem!

The following settings may be changed:

• Baud This cycle gadget may be used to set the baud rate of the serial port. As the serial port is only used to send command sequences and not to transfer real data, the baud rate doesn’t really matter; however, it should not be higher than the baud rate, your modem supports (default: ‘1200’).

DFA — the address utility for the AMIGA • Init String

The string that you enter into this string gadget, is used to init the modem. If you have a modem that supports the "Hayes" command sequences, the init command is usually ‘ATZ’ (default: ‘ATZ~~~’).

• Hangup String The string that is to be used to hang up the line, i.e. to terminate the connection, may be entered here (default: ‘+++\rTH0\r’)

• Predial String This string is sent to the modem just before the phone number itself. Some examples for "Hayes" compatible modems:

ATDP Pulse dial (default)

ATDT Tone dial

ATDP#W Makes the modem wait for a dialtone after sending a number (#). So it is possible to call from a extension line. Just substitute # by the number you would usually enter to get a dialtone.

• Suff. Dial String The string that has to be sent just after the phone number may be entered here. Please notice that this string has to be at least a ‘\r’, as the dialing does not work otherwise (default: ‘\r’).

• Serial device Here you may enter the name of a serial device that shall be used for dialing. Usually this is the Commodore ‘serial.device’, which is the default. If you have an interface card with (an) additional serial interface(s), you can insert the new serial device here.

• Select device Instead of typing the name of the serial device you can comfortably select it from a file requester. The selected device is inserted into the Serial device field.

• Device unit You may enter the number of the serial device here. Usually this is ‘0’. However, if you have an interface card with more than one serial port, they are distinguishable by their numbers (‘0’ and ‘1’ for example).

• Retries If a line is busy, it is possible to tell DFA to retry; the number of retries may be entered here.

• Redial delay The time delay between two dial events.

• Use OwnDevUnit The ‘OwnDevUnit.library’ gives you the chance to use DFA for dialing, even if the serial port is permanently occupied by a program as the ‘getty’ for example. The requirement for this to work is that this program supports the ‘OwnDevUnit.library’ as well of course. If you don’t use such a program, you may deactivate the Use OwnDevUnit checkbox. If there is no ‘OwnDevUnit.library’ installed on your computer, it can’t be used of course.

DFA — the address utility for the AMIGA

The input you can do into Init String, Hangup String, PredialString and Suff. Dial

String, can be found in the manual of your modem. In addition to the strings listed there, DFA supports the following command sequences:

‘~’ (Tilde) Causes a delay of about one second. ‘\b’ (Backspace) Deletes the character to its left. ‘\f’ (Formfeed) Performs a form feed. ‘\n’ (Linefeed) Produces a line feed. ‘\r’ (Carriage return) Performs a carriage return. ‘\t’ (Tab) Inserts a Tab character. ‘\e’ (ESC) Inserts an Escape character.

9.10 Paths and Filenames

The file and pathnames used by DFA may be set here.

More comfortable than to enter the pathname into the string gadgets manually is to hit the button just behind the corresponding string gadget. If you do so, a file requester pops up and you may select the desired file or pathname.

• Address file This is the address file that is loaded automatically during the startup of DFA (default: ‘s:adr.file’). This filename will be used as well, if you select Save as default.

• External files The external files are saved in a directory of their own (default: ‘s:DFAExternals’).

• File editor To modify the external files, you need an ascii editor. The ed, which can usually be found in the ‘C:’ drawer of your workbench, is an appropriate tool for example (default: ‘ed’).

• File viewer To view the external files, you need a text viewer, as for example the more program that comes along with your Workbench programs (default: ‘more’).

DFA — the address utility for the AMIGA

9.11 Arexx Scripts

It is possible to bind one Arexx script to every function key (F1 through F10). If you don’t like to enter the name of the Arexx script (including its path) by hand, you may alternatively hit the button gadget just behind the corresponding string gadget (keyboard shortcut: SHIFT <digit>). If you do so, a file requester will pop up and the desired Arexx script can be selected.

The Arexx scripts you enter here can only be reached via the function keys from the main window of the DFA-Editor. In the ‘rexx’ drawer of the DFA distribution some example scripts can be found; all files suited for this purpose have the extension ‘.dfa’.

Hit Ok to accept the changes or Cancel to quit the window and discard the changes.

9.12 Sorting

The addresses can be sorted by (nearly) every field of an address. To manage this, you may enter up to four sort criteria here. The most important one is Sort criteria 1. If two addresses can’t be distinguished by this criteria alone (e.g. if there are identical names), the second sort criteria (Sort criteria 2) is used. If a distinction is still impossible, the third sort criteria is used and so on.

• Sort Here you may choose which sort criteria you would like to change.

• Fields You may select one address field from this listview for every sort criteria (‘Sort criteria

    1’ through ‘Sort criteria 4’).

• Clear sort criteria If you want to delete a sort criteria, please hit this button.

• Ok and Cancel Ok accepts the changes. Please notice that the addresses are not resorted automatically. To sort your addresses using the modified sort criteria, you have to explicetly select Sort in the DFA-Editor (or use the corresponding Arexx command). Cancel ignores the changes.

Chapter 10: Legal stuff Page 71

10 Legal stuff

10.1 Licence

1. This licence applies to the product called ‘DFA’, a package for the AMIGA computer, published by Dirk Federlein under the concepts of Shareware. The terms "Package", "DFA" and "Archive" refer to this product. The licensee is addressed as "you". This license does not apply to third party supplied programs, found e.g. in the ‘c’ or ‘libs’ drawer of the original distribution archive.

2. You may use DFA for a period of up to 21 days for evaluation. After that, you have to register.

3. This archive may be freely redistributed only without the keyfile, which you got from me when you have registered. Please remember that the keyfile contains your name to prevent you from giving it away to others.

4. You are not allowed to draw commercial profit out of distributing this package. You may charge a fee to recover distribution and copy costs. The fee for diskette distribution may not be more than the cost to obtain a public domain diskette from Fred Fish.

5. You are not allowed to manipulate any file of DFA and distribute it under any or no name. Compression with common archivers like ‘zoo’ or ‘lharc’ is allowed. You are only allowed to redistribute the whole original archive (without the keyfile however). Some files may be modified for your personal use:

• The Arexx scripts The example Arexx scripts which can be found in the ‘Rexx’ drawer of the original distribution may be modified and/or enhanced.

• The locale files If you adopt the catalog translation files (extension ‘.ct’), you will be able to translate all program texts of DFA into your native language.

If you do such changes, you may do so for your personal use; redistribution of modified files is not allowed. (Nevertheless the author would be happy, if you’d sent him your modified files, so that they may serve as a basis for an inclusion into future versions of DFA)

6. If developers of commercial software would like to include certain parts of DFA in their commercial products, they will have to contact the author of DFA first, to ask for permission.

7. You agree to cease distributing DFA if requested to do so by the author.

8. By copying, distributing and/or using DFA you indicate your acceptance of this license to do so, and all its terms and conditions.

9. Each time you redistribute this package, the recipient automatically receives a license from the original licensor to copy, distribute and/or use this package subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herin.

10. As always, Fred Fish has got explicit permission to include this package into his wonderful "Fish Disk" collection, no matter if it is on floppy or compact disk.

10.2 Warranty

This software is provided as is without warranty of any kind, either expressed or implied. By using DFA, you agree to accept the entire risk as to the quality and performance of this package.

In no event (unless required by applicable law or agreed to in writing) will any copyright holder, or any other party who may redistribute this package as permitted above, be liable to you for damages, including any general, special, incidental or consequential damages arising out of the use or inability to use this package (including but not limited to loss of data or data being rendered inaccurate or losses sustained by you or third parties), even if such holder or other party has been advised of the possibility of such damages.

Of course, DFA has been tested rather extensively before it was released. In case you find any bugs in DFA I please you to send me a bug report. See Chapter 13 [Bugreports], page 75.

Chapter 11: Registration Page 73

11 Registration

Since Version 2.1 of DFA has been finished, it is an product with the SASG quality brand.

The advantages, which grow out of this can be looked up within the registration/information program.

Information about registration sites, the SASG, the shareware fee, payment methods etc. can be found there as well.

The registration program can be started either seperately or directly out of the about requester of DFA.

Before you register, you have the opportunity to test DFA in great detail. However, when using the unregistered version you can’t save the preferences (this feature is disabled); all changes you perform using the DFA-Preferences program stay active only until you reboot your computer. The window positions are therefore not saved as well.

In spite of all this restrictions it should be possible to see if DFA is what you expect. As soon as you use DFA regularly, however not later than after 21 days, you have to register!

You’ll get your personal keyfile, a TEX manual (in German and English) and an AmigaGuide manual (German and English). All manual files are on the disk only and not printed!

You’ll find my address below, see Chapter 14 [Important Addresses], page 76.

12 Updates

I’m continuously working on DFA and I’ll publish small updates from time to time. These updates will be fed into the following electronic nets:

• ADS Filenet

• AmiNet FTP Server

If you have access to such a electronic net, you can get the latest version of DFA quickly and cheaply. Many electronic mailboxes supply the latest version of DFA as well, e.g.

Publisher’s Treff 63739 Aschaffenburg Tel.: Germany–(0)6021–93435

Publisher’s Treff SASG Support BBS - All SASG products available 63739 Aschaffenburg Tel.: Germany-(0)6021-93435 Fido: 2:244/6302.0

You may get the latest version of DFA off from that BBS via FileRequest. The magic for the FileRequest is DFA. Any other SASG product may be requested as well.

Incubus

97074 Würzburg

Tel.: Germany–(0)931–781464, –781465, –781467, –781468, –72695, –884703, –884726

Alternatively you may order your update directly from me. Please use the registration program for this purpose — as you’ve already done when you’ve registered.

Fred Fish will get DFA to include it in his library. However, it will last a few weeks until DFA finds its was on these floppy or compact disks respectively.

Chapter 13: Bug Reports Page 75

13 Bug Reports

Coming with version 2.0 I did many significant changes to DFA. Usually this means that compared to DFA V1.2x there will be some new bugs in it — in spite of a long beta testing phase! This fact can be understood more easily, if you know that the source code of DFA takes up about 1.2 MB up to now!

So I ask you to report any bugs you discover. If you have any ideas what should be improved in future versions of DFA, please feel free to write, (mail or email, doesn’t matter) giving me any suggestions.

Bug reports can be sent by ordinary or elecronic mail. The address can be found in Chapter 14 [Important Addresses], page 76.

14 Important Addresses

14.1 The author’s address

Home address:

Dirk Federlein Brunnengasse 11a 63741 Aschaffenburg Germany

Study address:

Dirk Federlein Am Sonnenberg 34 97078 Würzburg Germany

E–Mail: UUCP: dirkf@alcmy.franken.de or dirk@alcmy.adsp.sub.org
FIDO: 2:244/6302.12

E-mail reaches me always, no matter to which of the above addresses you write. If you prefer to send "normal" mail, you should send it to Aschaffenburg.

Bank address:

Bayerische Vereinsbank Aschaffenburg Sort code: 795 200 70 Accont: 160 6 140 Germany

14.2 Registration in Australia

Andre Lackmann PO Box 259 Newport NSW 2106 Australia

E–Mail: UUCP: 76711.710@compuserve.com

Chapter 14: Important Addresses Page 77

14.3 Registration in England

Paul Jewell 7 Fairfield Avenue Victoria Park Cardiff Great Britain (Wales)

E–Mail: UUCP: Paul_Jewell@savanna.exnet.com

Bankverbindung:

Lloyds Bank Sort code: 30-98-94 Account: 7574426

14.4 Registration in France

Etienne Schneider 1 impasse des Sans Boutons F-80000 Amiens France

E–Mail: UUCP: etienne.schneider@ramses.fdn.org

14.5 Registration in Germany

Christoph Reichert Am Hasenkopf 4 63739 Aschaffenburg Deutschland

E–Mail: UUCP: chris@ptreff.ab.mayn.sub.de

Bank address:

Sparkasse Aschaffenburg - Alzenau Sort code: 795 500 00 Account: 528 01 77 Germany

14.6 Registration in the USA

Robert Blayzor PO Box 807 Johnstown, NY 12095-0807 U.S.A.

E–Mail: UUCP: die@lsd.org Chapter 15: Internal Page 79

15 Internal

15.1 Locale Support

While DFA V1.2x supported seven different languages, version 2.0 of DFA comes along only with english and german language support. This restriction has several reasons:

1. The english and german version of the catalog files can be done by myself. Therefore I am no longer — especially when creating updates — forced to get updated locale files from many different parts of this world in time.

2. For version 2.0 of DFA I enlarged the locale files by a big amount. To translate these files is a lot of work that I don’t want to be done by other people than by myself.

3. The enormous effort to adopt the locale files (for updates), prevented me a long time from doing a major update at all. Please notice that all locale files have to be changed, even if I add only one single sentence or word! Now there is no longer such a reason (for me) to delay updates!

4. The number of registered users that prefer other languages than german or english is still quite low. This has been the reason not to put a french locale into DFA V1.2x, as in the beginning there was no registered french user at all. Furthermore — most people prefer the english program texts even if English isn’t their native language.

Since version 2.1 has been finished, the french program texts are available as well!

15.2 The CheckBirthday Program

Since DFA got its Arexx–Port, it comes along with an Arexx script that displays coming birthdays. This script had, besides several pros, one real con: It needed much time to process a (long) address list.

Out of this reason DFA now comes along with an addition tool that does the same job, but much faster.

It can be started out of the shell, with the command:

     CheckBirthday ADR=ADRESSFILE/A BEFORE=DAYSBEFORE/N AFTER=DAYSAFTER/N DELAY/N

The meaning of the different parameters:

ADDRESSFILE

The address database file

BEFORE Number of days before the birthday

AFTER Number of days after the birthday

DELAY Delay in seconds after the output of all birthdays

Starting from Workbench is possible as well! Please notice that you have to make sure that you have adjusted the tooltypes to your needs. These tooltypes have the same names as the shell parameter. Additionally the WINDOW tooltype is available that gives you the chance to set the output window position and size, like:

     WINDOW=CON:20/20/400/200/Geburtstage/AUTO/CLOSE/WAIT

Chapter 16: Needed Configuration Page 81

16 Needed Configuration

The minimum configuration for the usage of DFA consists of any AMIGA computer that runs at least AMIGA–OS 2.0 and 1 MB Ram. As the addresses are completely kept in Ram, as more you have as better it is. A harddisk is not needed, but recommended. If you run (at least) Workbench 2.1 on your system, you may change the native language if you like (use the Locale preferences program of the Workbench to do so). If you run version 3.x of the operating system, you can enjoy some additional (mostly optical) features of this OS version.

17 Hints and Tips

Question:

After I restarted the computer, DFA has "forgotten" all window positions I’d adjusted before — why?

Answer:

To save the window positions, you have to save the DFA preferences. This can be done for example by selecting the appropriate menu item (‘Settings->Save Preferences’) of the DFA-Editor. Please remember that the settings and as a consequence of this the window positions as well, can only be saved in the registered version of DFA (see Chapter 11 [Registration], page 73, for details).

Question:

After I entered another position for the application icon, it stayed on the same place as before.

Answer:

Click on the Show check box gadget in the Miscellaneous Window of the DFA-Preferences program (cf. Section 9.4 [Miscellaneous], page 60) twice. Now the application icon should appear on the desired position, as long as there is no other icon that occupies this position on the Workbench.

Question:

As soon as I want to modify anything, a requester appears which tells me that I need "write access" before I may change my addresses.

Answer:

Switch off the network mode (see Section 9.4 [Miscellaneous], page 60). If you need the network mode, select Write Access in the Settings menu of the DFA-Editor.

Question:

If I hit the HELP key, I get no online help at all — nothing happens!

Answer:

Make sure that the online help files (‘DFAEditor.guide’ and ‘DFAPrefs.guide’) exist and the corresponding tooltypes of the DFA-Editor and the DFA-Preferences program are set correctly (cf. Section 9.1 [Tooltypes], page 58).

Question:

I appended an old DFA file to the current address database, but the appended addresses are not displayed within the address listview!

Answer:

Make sure that you activated the correct group flags. It might be that the appended addresses don’t have any group flags set. If this is the case, you will have Chapter 17: Hints and Tips Page 83

to deactivate all group check boxes in the DFA-Editor main window, to see these addresses.

Question:

If I access my addresses via Arexx, DFA seems to find not all addresses — sometimes it even finds no addresses at all!

Answer:

Please notice that the Arexx commands access only the addresses that correspond to the current group selection! To make sure that all addresses can be reached by Arexx, you should either select all group flags in the DFA-Editor main window or use the Arexx command CHANGEGROUPS accordingly.

Question:

After I’ve entered a new address containing group flags that did not match the current group selection, it was nevertheless displayed in the DFA-Editor’s main window listview. Isn’t this behaviour a bit inconsequent?

Answer:

Yes, indeed. However, for many people it would be even more confusing, if the new address which they’ve just entered, would no be displayed.

Question:

After I’ve entered a new address, for the present it was displayed in the DFA-Editor’s main window corretly. After I’ve resized the window, however, the address has disappeared!

Answer:

This problem should be seen together with the previous question. After a new address is entered, it is shown in the listview, no matter if the group selection fits or not. As soon as the window is resized, however, the listview entries are reformatted and the group selection is noticed! This causes all entries to disappear that do not match the current group selection.

Question:

Although I have set the modem preferences as it is needed, my modem does not dial any telephone numbers. The phone number of the address to dial is displayed within the dial window, but the modem does nothing that looks like dialing!

Answer:

Under some circumstances your modem does not like the baud rate you set in the System Preferences (Serial). Decrease the baud rate (in the System Preferences!), until your modem dials correctly.

18 Thanks to...

Many people helped me to make DFA what it is today. Therefore I would like to thank…

… Robert Blayzor

For supplying the registration site in the USA. … Michael ’Mick’ Hohmann

Tested the user aspects of DFA. … Martin ’XEN’ Huttenloher

Painted the wonderful MagicWB icons (8 colors) If you like the icons coming with this distribution of DFA, take a look at the complete collection of MagicWB icons by Martin Huttenloher.

… Kai Iske My "oldest" beta tester. He already tested DFA before version 1.0 has been published! For version 2.0 of DFA he supplied me a fast sort routine. The RemindBirthday utility is done by him as well.

… Katja and Paul Jewell

For the proof reading of the english manual and the english registration site. … Andre Lackmann

For supplying the registration site in Australia. … Etienne Schneider

For the french locale files and the french registration site. … Stefan Stuntz

For many telephone chats and his understanding that DFA still doesn’t support his Magic

User Interface (MUI).

… all others, that are not explicitly listed here and have sent me bug reports and/or suggestions for improving DFA

Thanks a lot to all DFA users to be so patient while waiting for DFA V2.0 such a long time. I do really hope it is worth the time you’ve waited for it!

Page i

Table of Contents

1 Welcome! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Features. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

4 Quick Start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

5 The concept of DFA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

6 Usage of DFA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

7 The DFA-Server program . . . . . . . . . . . . . . . . . . . . . . . . 10

7.1 Tooltypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 7.2 CLI–Parameter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 7.3 Commodity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 7.4 Application–Icon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 7.5 Arexx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

7.5.1 Important — new features since version 1.2x . . . . . . . . . 12 7.5.2 Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 7.5.3 Format of the Address Field. . . . . . . . . . . . . . . . . . . . . . . . . . 15 7.5.4 Arexx commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

7.5.4.1 The ABOUT Command. . . . . . . . . . . . . . . . . . . . . 16 7.5.4.2 The APPEND Command . . . . . . . . . . . . . . . . . . . 17 7.5.4.3 The ATTEMPTLOCK Command . . . . . . . . . . . 17 7.5.4.4 The CHANGEGROUPS Command . . . . . . . . . 18 7.5.4.5 The CLEARALL Command . . . . . . . . . . . . . . . . 19 7.5.4.6 The DESELECT Command. . . . . . . . . . . . . . . . . 19 7.5.4.7 The DIAL Command . . . . . . . . . . . . . . . . . . . . . . . 19 7.5.4.8 The EDIT Command . . . . . . . . . . . . . . . . . . . . . . . 20 7.5.4.9 The FIRST Command . . . . . . . . . . . . . . . . . . . . . . 22 7.5.4.10 The FREELOCK Command . . . . . . . . . . . . . . . 23 7.5.4.11 The GETCURRENT Command . . . . . . . . . . . 23 7.5.4.12 The GUI Command . . . . . . . . . . . . . . . . . . . . . . . 24 7.5.4.13 The ICONIFY Command . . . . . . . . . . . . . . . . . . 24 7.5.4.14 The KILL Command . . . . . . . . . . . . . . . . . . . . . . 25 7.5.4.15 The LAST Command . . . . . . . . . . . . . . . . . . . . . . 25 7.5.4.16 The LOAD Command . . . . . . . . . . . . . . . . . . . . . 26 7.5.4.17 The NEW Command . . . . . . . . . . . . . . . . . . . . . . 26 7.5.4.18 The NEWFILE Command . . . . . . . . . . . . . . . . . 28 Page ii DFA V2.1

7.5.4.19 The NEXT Command . . . . . . . . . . . . . . . . . . . . . 29 7.5.4.20 The NEXTSEL Command . . . . . . . . . . . . . . . . . 29 7.5.4.21 The OPENEDITWINDOW Command . . . . . 30 7.5.4.22 The POPUP Command . . . . . . . . . . . . . . . . . . . . 30 7.5.4.23 The PREV Command . . . . . . . . . . . . . . . . . . . . . 30 7.5.4.24 The PREVSEL Command . . . . . . . . . . . . . . . . . 31 7.5.4.25 The PRINT Command . . . . . . . . . . . . . . . . . . . . 31 7.5.4.26 The SAVE Command . . . . . . . . . . . . . . . . . . . . . . 32 7.5.4.27 The SAVEAS Command . . . . . . . . . . . . . . . . . . . 33 7.5.4.28 The SAVEASDEFAULT Command . . . . . . . . 33 7.5.4.29 The SEARCH Command . . . . . . . . . . . . . . . . . . 34 7.5.4.30 The SEARCHNEXT Command . . . . . . . . . . . . 35 7.5.4.31 The SELALL Command . . . . . . . . . . . . . . . . . . . 36 7.5.4.32 The SELBYNAME Command . . . . . . . . . . . . . 36 7.5.4.33 The SELECT Command . . . . . . . . . . . . . . . . . . . 37 7.5.4.34 The SORT Command. . . . . . . . . . . . . . . . . . . . . . 37

7.5.5 Error Codes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

8 The DFA-Editor program . . . . . . . . . . . . . . . . . . . . . . . . 41

8.1 Tooltypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 8.2 CLI parameter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 8.3 The Main Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 8.4 Tool Gadgets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 8.5 The Menu Items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

8.5.1 About . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 8.5.2 Info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 8.5.3 New file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 8.5.4 Open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 8.5.5 Append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 8.5.6 Save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 8.5.7 Save as. . .. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 8.5.8 Save as default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 8.5.9 Quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 8.5.10 Remove. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 8.5.11 Full. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 8.5.12 Edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 8.5.13 New . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 8.5.14 Kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 8.5.15 Copy to Clipboard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 8.5.16 (Un)set Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 8.5.17 Search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 8.5.18 Search Next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 8.5.19 Print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 8.5.20 Sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 8.5.21 Dial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 8.5.22 Preferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

Page iii

8.5.23 Edit Template . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
8.5.24 Write Access. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
8.5.25 Save Preferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

8.6 Full . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 8.7 Edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 8.8 New . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 8.9 Search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 8.10 Print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 8.11 Dial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

9 The DFA-Preferences Program . . . . . . . . . . . . . . . . . . 58

9.1 Tooltypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 9.2 CLI–Parameter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 9.3 The DFA-Preferences Main Window . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 9.4 Miscelleneous . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

9.4.1 Application Icon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 9.4.2 Auto Save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 9.4.3 Icon Creation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 9.4.4 Secret Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 9.4.5 Networking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

9.5 Address Listview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 9.6 Panel Listview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 9.7 Groups. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 9.8 Fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 9.9 Modem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 9.10 Paths and Filenames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 9.11 Arexx Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 9.12 Sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

10 Legal stuff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

10.1 Licence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 10.2 Warranty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

11 Registration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

12 Updates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

13 Bug Reports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

14 Important Addresses . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

14.1 The author’s address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 14.2 Registration in Australia . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 14.3 Registration in England . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 14.4 Registration in France. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 14.5 Registration in Germany . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 Page iv DFA V2.1

14.6 Registration in the USA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

15 Internal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

15.1 Locale Support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 15.2 The CheckBirthday Program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

16 Needed Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 17 Hints and Tips . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 18 Thanks to... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

Index

Index

(

(Un)set Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

A

About. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 59 Address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 Address file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 address file format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 address files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Address list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Address listview. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Address Listview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Address of the author . . . . . . . . . . . . . . . . . . . . . . . . 45, 59 Addresses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 ADDRESSFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 11 Adress of the author . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 All. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 AMIGA pattern. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 AppIcon. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44, 48 Application icon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Application Icon. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 60 Application Icon Position . . . . . . . . . . . . . . . . . . . . . . . . 60 Application Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Application–Icon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Arexx. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Arexx Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Arexx commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Arexx documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Arexx error codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Arexx Port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Arexx port name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Arexx return value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Arexx scripts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Arexx Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 ASCII editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 ASCII Viewer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Author’s address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Auto Save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

B

backup. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Bank address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76, 77, 78 bank transfer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76, 77, 78 Baud . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 Bug Reports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Button gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

C

Calling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 50, 68 Cancel (Edit) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Cancel (Print) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Cancel (Search) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Change Templates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 Check Box . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 CheckBirthday . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 Clear Column . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Clear line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Clear mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Clear sort criteria . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 CLI parameter of the DFA-Editor program . . . . . . . 42 CLI–Parameter des DFA-Preferences–Programms

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 CLI–Parameter des DFA-Servers . . . . . . . . . . . . . . . . . 11 Column . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Commodity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Copy to Clipboard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 59 CX POPKEY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 11 CX POPUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 CX PRIORITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Cycle Gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

D

default settings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Device unit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 DFAEDITOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 DFAPREFS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 42 Dial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 50, 57 Dialing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Documentation for using Arexx . . . . . . . . . . . . . . . . . . 13 DONOTWAIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

E

Edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 48, 52, 53 Edit Template . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 EMail address of the author . . . . . . . . . . . . . . . . . . . . . . 45 Error codes of Arexx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Exchange program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 External . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 External file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 External files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

F

Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 64, 70 Fields to search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 File editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 file format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 File Viewer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Filename. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 find . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Find . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 First . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 Format of the Address Field. . . . . . . . . . . . . . . . . . . . . . 15 format of the address file . . . . . . . . . . . . . . . . . . . . . . . . . 46 Formfeed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Full. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44, 48, 51

G

Generate Icon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 Giving a filename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Group handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65, 66 GUIDEFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 42, 58, 59

H

Hangup String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 Hayes compatible . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Hints and Tips . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 hotkey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 12 Hotkey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44, 48

I

Ignore case. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Important addresses. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 Info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Init String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Installer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Installieren . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 interface card . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Internal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

K

keyboard usage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 49

L

language . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Last . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Legal stuff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 Licence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 Lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Listview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Listview (main window) . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Load. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Locale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 locale support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

M

Mark all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Mark by name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Marking addresses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Memory Condition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Miscelleneous . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 Modem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

N

Native languages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 Needed Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 network mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 Network mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Networking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 New . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 48, 54 New address file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 New file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 New since version 1.2x . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Next Sel.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 None. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

O

Ok (Edit) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Ok (Full) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Online Help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 order of the address fields . . . . . . . . . . . . . . . . . . . . . . . . 13 OwnDevUnit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

P

Panel Listview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Panel–Listview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 Paths and Filenames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Phone. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 50, 57, 68 Port name for Arexx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 PORTNAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Position of the Application Icon . . . . . . . . . . . . . . . . . . 60 Predial String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Preferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 presets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Presets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 Prev. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Prev. Sel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Print. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 50, 56, 57 Print form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Print style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Print to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Public Screen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 58 PUBSCREEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 42, 58, 59

Q

Quick Start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Quit (DFA-Editor main window) . . . . . . . . . . . . . . . . . 44 Quit Preferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

R

Radio button. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Receiving Arexx commands . . . . . . . . . . . . . . . . . . . . . . 13 Redial delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Registration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Registration in Australia . . . . . . . . . . . . . . . . . . . . . . . . . 76 Registration in England . . . . . . . . . . . . . . . . . . . . . . . . . . 77 Registration in France. . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 Registration in Germany . . . . . . . . . . . . . . . . . . . . . . . . . 77 Registration in the USA. . . . . . . . . . . . . . . . . . . . . . . . . . 78 Remove . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Remove (DFA-Editor main window) . . . . . . . . . . 44, 48 Remove DFA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Retries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 68 Retry delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Return of address items . . . . . . . . . . . . . . . . . . . . . . . . . . 15 RXERR ȦPPENDFAILED . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ĊRTEXTIO. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ĊRTPORT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ĖNVCHANGED . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṀODIFIED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOCURRENT . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄODIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOEXTERNAL. . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOFILENAME . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOFIRST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOICONIFY . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOLAST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOLOAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOMEM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄONETWORKAPPEND . . . . . . . . . . . . . . . 39 RXERR ṄONETWORKCLEAR . . . . . . . . . . . . . . . . . 39 RXERR ṄONETWORKEDIT . . . . . . . . . . . . . . . . . . . 39 RXERR ṄONETWORKKILL . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄONETWORKNEW. . . . . . . . . . . . . . . . . . . . 39 RXERR ṄONETWORKSAVE . . . . . . . . . . . . . . . . . . . 39 RXERR ṄONETWORKSELECT . . . . . . . . . . . . . . . . 39 RXERR ṄONETWORKSORT . . . . . . . . . . . . . . . . . . . 39 RXERR ṄONEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄONEXTSEL. . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOPHONENUM . . . . . . . . . . . . . . . . . . . . . . . . 40 RXERR ṄOPREV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOPREVSEL . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄORELOAD. . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOSAVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOSERIALDEV . . . . . . . . . . . . . . . . . . . . . . . . 40 RXERR ṄOTFOUND . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṄOWRITELOCK . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ȮPNPRT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ṠERBUSY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 RXERR ṠERCONNECT . . . . . . . . . . . . . . . . . . . . . . . . . 40 RXERR ṠERERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 RXERR ṠERNOCARRIER . . . . . . . . . . . . . . . . . . . . . . 40 RXERR ṠERNODIALTONE . . . . . . . . . . . . . . . . . . . . . 40 RXERR ṠEROK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 RXERR ṠERRING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 RXERR ṠERUNKNOWN . . . . . . . . . . . . . . . . . . . . . . . . 40 RXERR ṠERVOICE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 RXERR ṠYNTAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ẆRITEALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 RXERR ẆRITECURRENT . . . . . . . . . . . . . . . . . . . . . 39 RXERR ẆRITESEL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

S

Save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 Save as default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 Save as. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 Save Preferences. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 59 search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 49, 54, 56 search criteria . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Search criteria . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Search for. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Search Next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 49 Secret Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 Select device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Selection: All . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Selection: By name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Selection: Clear . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Selection: Jump to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Serial device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 serial interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Serial number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 59 Serielles Gerät . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 shared access. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 shortcut. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Show Application Icon . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 Sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 50, 70 Sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 String Gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Suff. Dial String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

T

Templates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50, 54 Text Gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Text Viewer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Thanks to.... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 The CheckBirthday Program . . . . . . . . . . . . . . . . . . . . . 79 The concept of DFA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 The DFA-Editor program . . . . . . . . . . . . . . . . . . . . . . . . 41 The DFA-Preferences Main Window . . . . . . . . . . . . . . 59 The DFA-Server program . . . . . . . . . . . . . . . . . . . . . . . . 10 The Main Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 The Menu Items of the DFA-Editor . . . . . . . . . . . . . . 45 Then DFA-Preferences Program . . . . . . . . . . . . . . . . . . 58 Title . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 66 Tool Gadgets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Tooltypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Tooltypes for the DFA-Editor . . . . . . . . . . . . . . . . . . . . 41 Tooltypes of the DFA-Preferences programms. . . . . 58

U

Updates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 UPPERPUBSCREEN. . . . . . . . . . . . . . . . . . . . . . . 41, 42, 58, 59 Usage of DFA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Use Preferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Used . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

V

Verbergen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 version number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Version number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Visible lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

W

Warranty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Weight. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Write Access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50